### v3

# Accelerated Parallel Tempering via Neural Transports

Abstract
Markov Chain Monte Carlo (MCMC) algorithms are essential tools in computational statistics for sampling from unnormalised probability distributions, but can be fragile when targeting high-dimensional, multimodal, or complex target distributions. Parallel Tempering (PT) enhances MCMC’s sample efficiency through annealing and parallel computation, propagating samples from tractable reference distributions to intractable targets via state swapping across interpolating distributions. The effectiveness of PT is limited by the often minimal overlap between adjacent distributions in challenging problems, which requires increasing the computational resources to compensate. We introduce a framework that accelerates PT by leveraging neural samplers—including normalising flows, diffusion models, and controlled diffusions—to reduce the required overlap. Our approach utilises neural samplers in parallel, circumventing the computational burden of neural samplers while preserving the asymptotic consistency of classical PT. We demonstrate theoretically and empirically on a variety of multimodal sampling problems that our method improves sample quality, reduces the computational cost compared to classical PT, and enables efficient free energy/normalising constant estimation.
1 Introduction
Sampling from a probability distribution $\pi(x)=\exp(-U(x))/Z$ defined over a state-space $\mathcal{X}$ with tractable potential $U:\mathcal{X}\to\mathbb{R}$ and intractable normalising constant $Z=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$ is a fundamental task in machine learning and natural sciences. Markov Chain Monte Carlo (MCMC) methods are usually employed for such purposes, constructing an ergodic Markov chain $(X_{t})_{t\in\mathbb{N}}$ using local moves leaving $\pi$ invariant. While MCMC algorithms are guaranteed to converge asymptotically, in practice, they struggle when the target is complex with multiple well-separated modes (Papamarkou et al., 2022; Hénin et al., 2022). To handle such cases, Parallel Tempering (PT) (Swendsen & Wang, 1986; Geyer, 1991; Hukushima & Nemoto, 1996) is a popular class of MCMC methods designed to improve the global mixing of locally efficient MCMC algorithms.
PT works by considering an annealing path $\pi^{0},\pi^{1},\dots,\pi^{N}$ of distributions over $\mathcal{X}$ interpolating between a simple reference distribution $\pi^{0}=\eta$ (e.g. a Gaussian) and the target $\pi^{N}=\pi$. PT algorithms construct a Markov chain $\textbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$, targeting the joint distribution $\bm{\pi}=\pi^{0}\otimes\cdots\otimes\pi^{N}$. The PT chain $\mathbf{X}_{t}$ is constructed by alternating between (1) a local exploration phase where the $n$-th chain  [^1] Following the PT literature, we also refer to components of $\mathbf{X}_{t}$ as chains. of $\mathbf{X}_{t}$ is updated according to a $\pi^{n}$-invariant MCMC algorithm; and (2) a communication phase which proposes a sequence of swaps between neighbouring states accepted according to a Metropolis–Hastings correction ensuring invariance (see Figure˜1 (a)). Crucially, PT offsets the additional computation burden of simulating the extended $N$ chains through parallel computation, allowing for a similar effective computational cost as a single chain when implemented in a maximally parallelised manner.
Typically, the chains $X_{t}^{n}$ mix faster when closer to the reference and struggle closer towards the target. Therefore, communication between the reference and target, facilitated through swaps, can induce rapid mixing between modes of the target component (Woodard et al., 2009; Surjanovic et al., 2024). While the importance of the swapping mechanism for PT has led to a literature dedicated to optimising communication between the reference and target (Syed et al., 2021; 2022; Surjanovic et al., 2022) such works still rely on the original swapping mechanism (Geyer, 1991).
Figure 1: (Left) An illustration of the local exploration and communication step for PT vs APT. (Middle) 1,000 samples of a Gaussian mixture model target obtained using PT vs APT with a standard Gaussian reference. See Section˜6.1 for more details. (Right) Round trips for PT and APT with $N=6$ chains over $T=100,000$ iterations of Algorithm˜1.
As an alternative to PT, recent work has explored both continuous (Zhang & Chen, 2022; Vargas et al., 2023; Berner et al., 2024; Akhound-Sadegh et al., 2024; Vargas et al., 2024; Máté & Fleuret, 2023; Albergo & Vanden-Eijnden, 2024; Erives et al., 2025) and discrete (Noé et al., 2019; Papamakarios et al., 2021; Midgley et al., 2023; Gabrié et al., 2022) flows for sampling, under the umbrella of neural samplers. However, these methods usually incur a bias, foregoing theoretical guarantees of MCMC, and can be expensive to implement and train. Due to these shortcomings, standard PT provides a strong baseline that neural samplers struggle to match (He et al., 2025b).
Recent work (Arbel et al., 2021; Albergo & Vanden-Eijnden, 2024; Phillips et al., 2024; Chen et al., 2025) has explored approaches to debias neural samplers using Sequential Monte Carlo (SMC)-based ideas (Del Moral et al., 2006). Despite their consistency guarantees, these approaches do not address the mode-collapsing nature of many modern neural samplers (He et al., 2025b). PT offers a comparable but computationally dual framework to SMC, where the roles of parallelism and time are reversed (Syed et al., 2024). In SMC, particles are generated in parallel, and approximate annealing distributions are constructed sequentially. In contrast, in PT, particles are generated sequentially and annealing distributions are built in parallel. This raises the question: just as neural samplers have been integrated with SMC, can we integrate neural samplers with PT, combining the consistency of PT with the flexibility of neural samplers?
We answer positively this question by formalising and exploiting the framework introduced by Ballard & Jarzynski (2009; 2012) in physics for designing more flexible swap mechanisms, which we call accelerated PT (APT). APT preserves PT’s asymptotic consistency and allows us to integrate easily normalising flows, stochastic control, and diffusions into existing PT implementations. ATP uses neural sampling in a parallelised manner, mitigating its high computational burden. Empirically, it outperforms PT by accelerating the communication between the reference and target states, even when controlling for the additional computation incurred by neural samplers.
A relevant prior work integrating PT with normalising flows is Invernizzi et al. (2022): it trains a normalising flow to directly map configurations from the highest-temperature reference distribution of a molecular system to the lowest-temperature target distribution, effectively bypassing the intermediate annealing distributions. By contrast, our framework leverages normalising flows to facilitate exchanges between all neighbouring temperature levels, thereby enhancing sampling efficiency across the entire annealing path and providing a more stable training objective.
2 Parallel Tempering
Let $\pi^{0},\pi^{1},\dots,\pi^{N}$ be an annealing path of probability distributions on $\mathcal{X}$ where $\pi^{0}=\eta$ is the reference and $\pi^{N}=\pi$ is the target. We assume the $n$-th annealing distribution admits density $\pi^{n}(x):=\exp(-U^{n}(x))/Z_{n}$ with respect to a base measure $\mathrm{d}x$, where $U^{n}:\mathcal{X}\to\mathbb{R}$ is the potential and $Z_{n}:=\int_{\mathcal{X}}\exp(-U(x))\mathrm{d}x$. We further assume we can evaluate the work $W^{n}:\mathcal{X}\to\mathbb{R}$ between $\pi^{n-1}$ and $\pi^{n}$, equal to the change in log-likelihood of the un-normalised annealing distributions:
$W^{n}(x):=\Delta F_{n}-\log\frac{\mathrm{d}\pi^{n}}{\mathrm{d}\pi^{n-1}}(x)=U^{n}(x)-U^{n-1}(x),$ (1)
where $\Delta F_{n}:=\log Z_{n-1}-\log Z_{n}$. Our goal is to estimate $\pi[f]:=\int_{\mathcal{X}}f(x)\pi(\mathrm{d}x)$, the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi$, and the change in free energy $\Delta F:=\log Z_{0}-\log Z$.
There is considerable flexibility in choosing annealing distributions provided $U^{0}(x)=\log\eta(x)$ with $Z_{0}=1$ and $U^{N}(x)=U(x)$ with $Z_{N}=Z$. Without loss of generality, we can assume $U^{n}=U^{\beta_{n}}$ where for $\beta\in[0,1]$, $U^{\beta}:\mathcal{X}\to\mathbb{R}$ continuously interpolates between reference and target potentials as $\beta$ increases from 0 to $1$, according to the annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$. A common choice is the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. See Masrani et al. (2021); Syed et al. (2021); Máté & Fleuret (2023); York (2023) for alternative non-linear annealing paths.
2.1 Non-Reversible Parallel Tempering
The PT algorithm constructs a Markov chain $\mathbf{X}_{t}=(X_{t}^{0},\ldots,X_{t}^{N})$ on the extended state-space $\mathcal{X}^{N+1}$ invariant to the joint distribution $\bm{\pi}:=\pi^{0}\otimes\cdots\otimes\pi^{N}$. We construct $\mathbf{X}_{t}$ from $\mathbf{X}_{t-1}$ by doing (1) a local exploration step followed by (2) a communication step seen in the top of Figure˜1 (a). For $n=0,\dots,N$, the $n$-th local exploration move (1) updates the $n$-th component of $\mathbf{X}_{t-1}$ using a $\pi^{n}$-invariant Markov kernel $K^{n}(x,\mathrm{d}x^{\prime})$ on $\mathcal{X}$ corresponding to an MCMC move targetting $\pi^{n}$,
$X^{n}_{t}\sim K^{n}(X_{t-1}^{n},\mathrm{d}x^{n}),$
We additionally assume that $K^{0}(x,\mathrm{d}x^{\prime})=\eta(\mathrm{d}x^{\prime})$ corresponds to an independent sample from the reference. The communication step (2) applies a sequence of swap moves between adjacent components of $\mathbf{X}_{t}$, where the $n$-th swap move illustrated in Figure˜1 exchanges components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ with probability $\alpha^{n}(X^{n-1}_{t},X^{n}_{t})$, where for $x,x^{\prime}\in\mathcal{X}$,
$\alpha^{n}(x,x^{\prime}):=\exp\left(\min\left\{0,W^{n}(x^{\prime})-W^{n}(x)\right\}\right).$ (2)
In practice, it is advantageous to use a non-reversible communication (Okabe et al., 2001; Syed et al., 2022), where the $n$-th swap move is proposed only at iterations with matching parity, $n\equiv t\mod 2$. Both local exploration and communication steps can be done in parallel, allowing distributed implementations to leverage parallel computation to accelerate sampling (Surjanovic et al., 2023).
Round trips
While the effective sample size (ESS) of samples generated by a Markov chain is the gold standard for evaluating the performance of MCMC algorithms, in our setting, we are mainly interested in improving the swap kernel within PT. As ESS measures the intertwined performance of the local exploration and swap kernels, we are instead interested in maximising communication between reference and target, which can be empirically measured by counting the total number of round trips, $R_{T}$ tracking the number of independent reference samples transported to the target after $T$ iterations of PT (Katzgraber et al., 2006; Lingenheil et al., 2009). See Section˜A.1 for a formal definition. The mixing time of the PT chain is related to the time it takes for a round trip to occur and the total number of round trips is a measure of the particle diversity generated by PT and strongly correlates with the ESS (Surjanovic et al., 2024).
Algorithm 1 Accelerated Parallel Tempering
1:Initialise $\textbf{X}_{0}=(X^{0}_{0},\dots,X^{N}_{0})$;
2:for $t=1,\dots,T$ do
3:  $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t}),\quad X^{n}_{t}\sim K^{n}(X^{n}_{t-1},\mathrm{d}x)\triangleright$ Local exploration move
[…]
5:   $\accentset{\rightharpoonup}{X}^{n-1}_{t,0},\accentset{\leftharpoonup}{X}^{n}_{t,K}\leftarrow X^{n-1}_{t},X^{n}_{t}\triangleright$ Initialise forward/backward paths
6:   for $k=1,\dots,K$ do
7:     $\accentset{\rightharpoonup}{X}^{n-1}_{t,k}\sim P_{k}^{n-1}(\accentset{\rightharpoonup}{X}^{n-1}_{t,{k-1}},\mathrm{d}x)$ $\triangleright$ Accelerate forward
8:     $\accentset{\leftharpoonup}{X}^{n}_{t,K-k}\sim Q_{K-k}^{n}(\accentset{\leftharpoonup}{X}^{n}_{t,K-k+1},\mathrm{d}x)$ $\triangleright$ Accelerate backward
9:   end for
10:   $\accentset{\rightharpoonup}{W}^{n}_{K,t},\accentset{\leftharpoonup}{W}^{n}_{K,t}\leftarrow W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}),W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})\triangleright$ Work of forward/backward paths
11:   $U\sim\mathrm{Uniform}([0,1])$
12:   if $\log U<\accentset{\rightharpoonup}{W}^{n}_{K,t}-\accentset{\leftharpoonup}{W}^{n}_{K,t}$ then $\triangleright$ Accelerated swap move
13:     $X^{n-1}_{t},X^{n}_{t}\leftarrow X^{n}_{t,0},X^{n-1}_{t,K}$
14:   end if
15:  end for
16:end for
17:Return: $\textbf{X}_{1},\dots,\textbf{X}_{T}$
3 Accelerated Parallel Tempering
A limitation of PT is its inflexible swap move, which only proposes directly exchangeable samples between distributions, considering just the relative change in potentials $\pi^{n-1}$ and $\pi^{n}$. Low acceptance probability occurs when these distributions have minimal overlap. Addressing this requires increasing the number of parallel chains $N$, which may not always be possible. We propose Accelerated PT (APT), expanding the framework developed in Ballard & Jarzynski (2009; 2012) to improve distributional overlap and accelerate communication for a fixed number of chains.
3.1 Forward and Backward Accelerators
For $n=1,\dots,N$, let $P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ and $Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$, $k=1,\dots,K$, be two families of $K$ transition kernels which we call forward accelerators and backward accelerators. They induce ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$, two time-inhomogeneous Markov processes obtained by propagating $\pi^{n-1}$ forward in time and $\pi^{n}$ backward in time respectively using the forward and the backward accelerators,
${\mathbb{P}}^{n-1}_{K}(\mathrm{d}x_{0:K})\!:=\!\pi^{n-1}(\mathrm{d}x_{0})\!\prod_{k=1}^{K}\!P^{n-1}_{k}(x_{k-1},\mathrm{d}x_{k}),{\mathbb{Q}}^{n}_{K}(\mathrm{d}x_{0:K})\!:=\!\pi^{n}(\mathrm{d}x_{K})\!\prod_{k=1}^{K}\!Q^{n}_{k-1}(x_{k},\mathrm{d}x_{k-1}).$
We assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ are mutually absolutely continuous and we can evaluate $W^{n}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$, the work between the forward and the backward accelerated paths, extending $W^{n}$ in Equation˜1,
$W^{n}_{K}(x_{0:K}):=\Delta F_{n}-\log\frac{\mathrm{d}{{\mathbb{Q}}}^{n}_{K}}{\mathrm{d}{{\mathbb{P}}}^{n-1}_{K}}(x_{0:K}).$ (3)
3.2 Non-Reversible Accelerated Parallel Tempering
We construct the Markov chain $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})$ for $t=1,\dots,T$ using the same local exploration and non-reversible communication as classical PT, but we use using the accelerated PT swap move as described below and summarized in Algorithm˜1. Given PT state $\textbf{X}_{t}=(X^{0}_{t},\dots,X^{N}_{t})\in\mathcal{X}^{N+1}$ after the local exploration move, we define the $n$-th accelerated swap move as follows: generate paths $\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K}$, and $\accentset{\leftharpoonup}{X}^{n}_{t,0:K}$ obtained by propagating $X^{n-1}_{t}$ and $X^{n}_{t}$ forward and backward in time using the forward and backward transitions respectively,
$\accentset{\rightharpoonup}{X}^{n-1}_{t,0}=X^{n-1}_{t},\quad\accentset{\rightharpoonup}{X}_{t,k}^{n-1}\sim P^{n-1}_{k}(\accentset{\rightharpoonup}{X}^{n-1}_{t,k-1},\mathrm{d}x_{k}),$
$\accentset{\leftharpoonup}{X}^{n}_{t,K}=X^{n}_{t},\quad\accentset{\leftharpoonup}{X}_{t,k-1}^{n}\sim Q^{n}_{k-1}(\accentset{\leftharpoonup}{X}^{n}_{t,k},\mathrm{d}x_{k-1}).$
The $n$-th accelerated swap move illustrated in Figure˜1 replaces components $X^{n-1}_{t}$ and $X^{n}_{t}$ in $\textbf{X}_{t}$ with $\accentset{\leftharpoonup}{X}^{n}_{t,0}$ and $\accentset{\rightharpoonup}{X}^{n-1}_{t,K}$ respectively with probability $\alpha^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K},\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, where for $x_{0:K},x_{0:K}^{\prime}\in\mathcal{X}^{K+1}$,
$\alpha_{K}^{n}(x_{0:K},x_{0:K}^{\prime})=\exp\left(\min\left\{0,W_{K}^{n}(x_{0:K}^{\prime})-W_{K}^{n}(x_{0:K})\right\}\right).$
When $K=0$, the accelerated swap coincides with the traditional PT swap in Equation˜2. However, an accelerated swap can obtain an acceptance of $1$ even if $\pi^{n-1}\neq\pi^{n}$ provided ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$. Therefore, we aim to choose the forward and backward accelerators to make the laws of simulated forward and backward paths as close to each other as possible. Theorem˜1 shows we can quantify this discrepancy in through the rejection rate, $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}$ which by Pinsker inequality is controlled by the symmetric KL divergence between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$,
$r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})^{2}\leq\frac{1}{2}{\mathbb{P}}^{n-1}_{K}[-W^{n}_{K}]+\frac{1}{2}{\mathbb{Q}}^{n}_{K}[W^{n}_{K}]=:\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}).$ (4)
Theorem 1.
The APT Markov chain $\textbf{X}_{t}$ generated by Algorithm˜1 is ergodic and $\bm{\pi}$-invariant. Moreover, the probability the $n$-th accelerated swap is rejected at stationarity equals $r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$.
Expectation and free energy estimators
As a by-product of Algorithm˜1, we obtain a consistent estimator $\hat{\pi}^{n}_{T}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n})$ for the expectation of $f:\mathcal{X}\to\mathbb{R}$ with respect to $\pi^{n}$ by taking the Monte Carlo average over the $n$-th chain $X^{n}_{t}$. To also obtain free energy estimates, by averaging over $\accentset{\rightharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ and $\accentset{\leftharpoonup}{W}^{n}_{K,t}:=W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{t,0:K})$, the work done by the forward and backward accelerated paths generated during the $n$-th accelerated swap at time $t$ during the communication phase of Algorithm˜1, to obtain consistent estimators $\Delta\accentset{\rightharpoonup}{F}_{T}$ and $\Delta\accentset{\leftharpoonup}{F}_{T}$ for $\Delta F$ respectively as $T\to\infty$,
$\exp(-\Delta\accentset{\rightharpoonup}{F}_{T})\!:=\!\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\exp\left(-\accentset{\rightharpoonup}{W}^{n}_{K,t}\right),~\exp(\Delta\accentset{\leftharpoonup}{F}_{T})\!:=\!\prod_{n=1}^{N}\frac{2}{T}\sum_{n\equiv t\mod 2}\exp\left(\accentset{\leftharpoonup}{W}^{n}_{K,t}\right).$
We then obtain a consistent estimator $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ for $\Delta F$ by averaging the forward and backward estimators. When we consider the classical PT with $K=0$, these estimators degrade back to free energy perturbation (FEP) (Zwanzig, 1954); when the path is deterministic (e.g., implemented through normalising flow), these estimators recover the target FEP (Jarzynski, 2002); and when the path is stochastic, the estimators can be viewed as a form of Jarzynski equality (Jarzynski, 1997) or escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). Additionally, following He et al. (2025a, Proposition 3.2), we can also apply Bennett acceptance ratio using the work for the paths (Bennett, 1976; Shirts et al., 2003; Hahn & Then, 2009; Minh & Chodera, 2009; Vaikuntanathan & Jarzynski, 2011) to achieve reduced variance without additional functional calls.
Proposition 1.
The estimators $\hat{\pi}^{n}_{T}[f]$ and $\Delta\hat{F}_{T}$ a.s. converge to $\pi^{n}[f]$ and $\Delta F$ respectively as $T\to\infty$. Moreover, if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$ for all $n$, then $\Delta\hat{F}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F$.
4 Analysis of Accelerated PT
We analyse how the performance, measured by the round trip observed after $T$ iterations, $R_{T}$, scales with increasing parallel chains $N$ and acceleration time $K$. It is equivalent to analyse the round trip rate $\tau:=\lim_{T\to\infty}\mathbb{E}[R_{T}]/T$, defined as the expected percentage of PT iterations, where a round trip occurs (Katzgraber et al., 2006; Lingenheil et al., 2009), which coincides with slope in Figure˜1 (Right). To derive theoretical insights independent of the problem-specific local exploration move, we make an efficient local exploration assumption analogous to Syed et al. (2021; 2022); Surjanovic et al. (2024), which assumes the work done by forward and backward accelerated paths is independent across chains and PT iterations.
Assumption 1 (Efficient local exploration).
For all $n=1,\dots,N$, $(\accentset{\rightharpoonup}{W}^{n}_{K,t},\accentset{\leftharpoonup}{W}^{n}_{K,t})$ are iid in $t$ and equal in distribution to $(W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{0:K}),W^{n}_{K}(\accentset{\leftharpoonup}{X}^{n}_{0:K}))$ where $(\accentset{\rightharpoonup}{X}^{n-1}_{0:K},\accentset{\leftharpoonup}{X}^{n}_{0:K})\sim{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}$.
Proposition˜2 relates the round trip rate to the rejections, extending Syed et al. (2021, Corollary 2).
Proposition 2.
If ˜1 holds, then $\tau=\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ where,
$\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K}):=\left(2+2\sum_{n=1}^{N}\frac{r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})}{1-r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})}\right)^{-1}.$
Asymptotic scaling with acceleration time.
We now fix $N$ and use Proposition˜2 to study how the round trip rate scales with $K$. We focus on a special case where ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ arise as $K$-step discretisations of an underlying stochastic differential equation (SDE) which bridge $\pi^{n-1}$ and $\pi^{n}$ with path measure ${\mathbb{P}}^{n-1}_{\infty}$ and ${\mathbb{Q}}^{n}_{\infty}$ as $K\to\infty$.
Proposition 3.
Under appropriate conditions on the drifts of the SDE (Section˜B.2), as $K\to\infty$, $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\tau({\mathbb{P}}^{0:N-1}_{\infty},{\mathbb{Q}}^{1:N}_{\infty})$ and $r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})+\mathcal{O}(\frac{1}{\sqrt{K}})$.
The main appeal of accelerated PT comes from the fact that for good neural network approximations
$r({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty})\approx 0\ll r(\pi^{n-1},\pi^{n}).$
Thus, increasing $K$ can dramatically reduce the rejection rate and increase the round trip rate in challenging scenarios. In practice, we find that for moderate values of $K$, this benefit overweights the increased computational cost as seen in Table˜1.
Asymptotic scaling with increased parallelism.
To understand how accelerated PT scales with increased parallelism, as $N\to\infty$ for a fixed number of acceleration steps $K$. For each $N$, we assume $\pi^{n}$ discretizes a continuous annealing path $\pi^{\beta}$ interpolating between reference and target at $\beta=\beta_{n}$ for some annealing schedule $0=\beta_{0}<\dots<\beta_{n}=1$. Additionally we assume ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ discretize the mutually absolutely continuous forward/backward accelerators ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ between $\pi_{\beta}$ and $\pi_{\beta^{\prime}}$ at $\beta=\beta_{n-1}$ and $\beta^{\prime}=\beta_{n}$ respectively with work function $W^{\beta,\beta^{\prime}}_{K}$.
Theorem 2.
Suppose ${\mathbb{P}}^{\beta,\beta^{\prime}}_{K}$ and ${\mathbb{Q}}^{\beta,\beta^{\prime}}_{K}$ are sufficiently regular and satisfy Assumptions 2–4 in Section˜B.3. As $N\to\infty$ if $\max_{n\leq N}|\beta_{n}-\beta_{n-1}|=O(N^{-1})$, then $\sum_{n=1}^{N}r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$ converges to $\Lambda_{K}$ and $\tau({\mathbb{P}}^{0:N-1}_{K},{\mathbb{Q}}^{1:N}_{K})$ converges to $\bar{\tau}_{K}=(2+2\Lambda_{K})^{-1}$, where $\Lambda_{K}$ equals,
$\Lambda_{K}:=\int_{0}^{1}\frac{1}{2}\mathbb{E}[|\dot{W}^{\beta}_{K}(\accentset{\leftharpoonup}{X}^{\beta}_{0:K})-\dot{W}^{\beta}_{K}(\accentset{\rightharpoonup}{X}^{\beta}_{0:K})|]\mathrm{d}\beta,\quad(\accentset{\rightharpoonup}{X}_{0:K}^{\beta}\accentset{\leftharpoonup}{X}^{\beta}_{0:K})\sim{\mathbb{P}}^{\beta,\beta}_{K}\otimes{\mathbb{Q}}^{\beta,\beta}_{K},$
and $\dot{W}^{\beta}_{K}:\mathcal{X}^{K+1}\to\mathbb{R}$ is the partial derivative with respect to $\beta^{\prime}$ of $W^{\beta,\beta^{\prime}}_{K}$ at $\beta^{\prime}=\beta$.
Theorem˜2 shows that as $N\to\infty$, the round trip rate is controlled by the global barrier for APT, $\Lambda_{K}$, and can be approximated by rejection rates obtained in Algorithm 1. Notably, APT observes a sharp deterioration in round trip when $N\ll\Lambda_{K}$, begins to stabilise when $N\approx\Lambda_{K}$ and observes a marginal improvement when $N\gg\Lambda_{K}$. As discussed in Proposition˜3, for a reasonable choice of accelerators we expect $\Lambda_{K}$ to decrease with $K$, which suggests APT can substantially improve compared to classical PT ($K=0$) for challenging problems where $N\approx\Lambda_{0}>\Lambda_{K}$. Finally, we explore the trade-off between increasing parallelism and increasing acceleration time in Appendix˜B.
5 Design Space for Diffusion PT
5.1 Normalising Flow Accelerated PT
Normalising flows (Tabak & Vanden-Eijnden, 2010; Rezende & Mohamed, 2015) provide a flexible framework for approximating complex probability distributions by transforming a simple base distribution (e.g., a Gaussian) through a sequence of invertible, differentiable mappings.
Let $T^{n}:\mathcal{X}\to\mathcal{X}$ be such a diffeomorphism. We denote NF-APT as APT with forward and backward accelerators to deterministically transport by normalising flow $T^{n}$ and its inverse, respectively, motivated by the use of normalising flows to map between annealing distributions in Arbel et al. (2021) and Matthews et al. (2022) in the context of SMC samplers (Del Moral et al., 2006),
$P_{1}^{n-1}(x_{0},\mathrm{d}x_{1})=\delta_{T^{n}(x_{0})}(\mathrm{d}x_{1}),\quad Q_{0}^{n}(x_{1},\mathrm{d}x_{0})=\delta_{(T^{n})^{-1}(x_{1})}(\mathrm{d}x_{0}).$
Then the work functional equals,
$W^{n}_{1}(x_{0},x_{1})=U^{n}(x_{1})-U^{n-1}(x_{0})-\log|\det\nabla T^{n}(x_{0})|,\quad x_{1}=T^{n}(x_{0}).$
We can parametrise the normalising flow $T^{n}$ by a neural network trained to optimise the symmetric KL-divergence from Equation˜4, $\mathcal{L}(T)=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})$. See Section˜C.1 for details.
5.2 Controlled Monte Carlo Diffusions Accelerated PT
Another common transport for sampling is based on the escorted Jarzynski equality (Vaikuntanathan & Jarzynski, 2008). One way to realise this in practice with neural networks is via Controlled Monte Carlo Diffusions (CMCD) (Vargas et al., 2024). Suppose $\mathcal{X}\subseteq\mathbb{R}^{d}$ and for $s\in[0,1]$ we have $b^{n}_{s}\in\mathbb{R}^{d}$, $\sigma^{n}_{s}\in\mathbb{R}_{+}$, and $U^{n}_{s}=(1-\phi_{s}^{n})U^{n-1}+\phi_{s}^{n}U^{n}$, where $\phi_{s}^{n}\in[0,1]$ is a monotonically increasing in $s$ and $\phi^{n}_{0}=0,\phi^{n}_{1}=1$. For a fixed $K$, let $s_{k}=k/K$ and $\Delta s_{k}=1/K$ be the uniform discretisation of the unit interval. We denote CMCD-APT as APT with $k$-th forward transition kernel,
$P^{n-1}_{k}(x_{k-1},\mathrm{d}x_{k})=\mathcal{N}(x_{k-1}-(\sigma^{n}_{s_{k-1}})^{2}\nabla U^{n}_{s_{k-1}}(x_{k-1})\Delta s_{k}+b^{n}_{s_{k-1}}(x_{k-1})\Delta s_{k},2(\sigma^{n}_{s_{k-1}})^{2}\Delta s_{k}),$
and the backward transition,
$Q^{n}_{k-1}(x_{k},\mathrm{d}x_{k-1})=\mathcal{N}(x_{k}+(\sigma_{s_{k}}^{n})^{2}\nabla U^{n}_{s_{k}}(x_{k})\Delta s_{k}+b^{n}_{s_{k}}(x_{k})\Delta s_{k},2(\sigma_{s_{k}}^{n})^{2}\Delta s_{k}).$
The corresponding work function for CMCD-APT equals,
$W^{n}_{K}(x_{0:K})=U^{n}(x_{K})-U^{n-1}(x_{0})+\sum_{k=1}^{K}\log P^{n-1}_{k}(x_{k-1},x_{k})-\sum_{k=1}^{K}\log Q^{n}_{k-1}(x_{k},x_{k-1}),$ (5)
where $P^{n-1}_{k}(x_{k-1},x_{k})$ and $Q^{n}_{k}(x_{k},x_{k-1})$ are the Gaussian densities of the forward and backwards kernel transitions with respect to the Lebesgue measure. We learn the neural transport $b^{n}_{s}$, the time schedule $\phi^{n}_{s}$ and the diffusion coefficient $\sigma^{n}_{s}$. We use a symmetrised KL-divergence loss $\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}^{n}_{K})$, where we also learn $\phi_{s}^{n},\sigma^{n}_{s}$ following Geffner & Domke (2023) [^2] We note that in standard CMCD, we cannot use symmetrised KL-divergence as the objective, as we only have access to data from one side.. While in theory, once the vectorised field $b_{s}^{n}$ is perfectly learned, any $\sigma^{n}_{s}$ can be used, we found that learning $\sigma^{n}_{s}$ can significantly stabilise the training and enhance the results. Note that other divergences and losses can be used, such as (Máté & Fleuret, 2023; Richter & Berner, 2023; Albergo & Vanden-Eijnden, 2024). We discuss the form of the transition kernels, the work, and loss in the limit at $K\to\infty$ in Section˜C.2.
5.3 Diffusion Accelerated PT
Given $\mathcal{X}=\mathbb{R}^{d}$, the Variance-Preserving (VP) diffusion model (Ho et al., 2020; Song et al., 2020) is defined by the SDE $dY_{s}=-\gamma_{s}Y_{s}\mathrm{d}s+\sqrt{2\gamma_{s}}\mathrm{d}{W}_{s}$, with $s\in[0,1]$, $Y_{0}\sim\pi$, and $\gamma_{s}$ is a rate function from $[0,1]$ to $\mathbb{R}^{+}$. The time-reversal SDE $(X_{s})_{s\in[0,1]}=(Y_{1-s})_{s\in[0,1]}$ has the form $dX_{s}=[\gamma_{1-s}X_{s}+2\gamma_{1-s}\nabla\log\pi_{s}^{\text{VP}}(X_{s})]\mathrm{d}s+\sqrt{2\gamma_{1-s}}\mathrm{d}{W}_{s}$ where $\pi_{s}^{\text{VP}}$ is the density of $Y_{1-s}$ and $X_{0}\sim\pi_{0}^{\text{VP}}={\mathcal{N}}(0,\mathrm{I})$ is close to a standard Gaussian in the common scenario that $\int_{0}^{1}\gamma_{s}\mathrm{d}s\gg 1$. The score $\nabla\log\pi_{s}^{\text{VP}}$ is unknown but can be learned by score-matching objectives.
We introduce for $s\in[0,1]$ an energy-based model (Salimans & Ho, 2021) with potential $\tilde{U}^{\text{VP}}_{s}:\mathcal{X}\mapsto\mathbb{R}$ and corresponding annealing distributions $\tilde{\pi}^{\text{VP}}_{s}\propto\exp(-\tilde{U}^{\text{VP}}_{s})$ to approximate $\pi_{s}^{\text{VP}}$ while satisfying the boundary conditions $\tilde{\pi}^{\text{VP}}_{0}={\mathcal{N}}(0,\mathrm{I})$ and $\tilde{\pi}^{\text{VP}}_{1}=\pi$. For some schedule $0=s_{0}<\cdots<s_{N}=1$, we consider the annealing path $\pi^{n}=\tilde{\pi}^{\text{VP}}_{s_{n}}$. For a fixed $K$, we denote Diff-APT as APT with forward and backward kernels respectively given by some $K$-step integrator of the time-reverse SDE and the forward SDE between $s_{n-1}$ and $s_{n}$. For instance, the Euler integrator with step-size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$ yields:
$P_{k}^{n-1}(x_{k-1},\mathrm{d}x_{k})$ $=\mathcal{N}(x_{k-1}+\delta_{n}\gamma_{1-s_{n,k}}x_{k-1}-2\gamma_{1-s_{n,k}}\nabla\tilde{U}_{s_{n,k},\theta}^{\text{VP}}(x_{k-1}),2\delta_{n})$ (6)
$Q_{k-1}^{n}(x_{k},\mathrm{d}x_{k-1})$ $=\mathcal{N}(x_{k}-\delta_{n}\gamma_{1-s_{n,k}}x_{k},2\delta_{n}).$ (7)
The resulting work function $W^{n}_{K}$ coincides with Equation˜5 from CMCD-APT, replacing the forward backward transitions $P^{n-1}_{k}$ and $Q^{n}_{n-1}$ for CMCD-APT with Equation˜6 and Equation˜7. Further, we parametrise $\tilde{U}_{s}^{\text{VP}}$ by a neural network trained to minimise the standard score matching objective. Since we cannot access samples from $X_{1}\sim\pi$, we iteratively train on approximate samples from $\pi$ obtained by sampling from Diff-APT. We stress that a parallelised implementation of Diff-APT allows for generating a sample from $\pi$ with a similar effective cost as a single discretisation step solving the time-reversal SDE. For further details, see Section˜C.3.
6 Experiments
In this section, we evaluate our proposed accelerated parallel tempering and its three variants on a variety of targets. We defer all experimental details to Appendix˜D.
6.1 Comparison of Acceleration Methods
We evaluate the three variants of APT introduced in Section˜5: Flow-APT, CMCD-APT and Diff-APT, against the baseline of PT using the linear path with reference $\pi_{0}={\mathcal{N}}(0,\mathrm{I})$, on a 40-mode Gaussian mixture model in 10 dimensions (GMM-$10$) (Midgley et al., 2023). Table˜1 compares the performance of PT vs the APT variants for $N=6,10,30$ chains. We choose these numbers to represent: (1) the smallest $N$ where standard PT obtains round trips (2) the regime $N\approx\Lambda$ where PT is stable, and (3) the regime where $N\gg\Lambda$ and PT is close to optimal. All three APT variants substantially improve over classical PT, with increased round trips with $N$ and $K$. In the small $N$ regime (1) resulted in a 10x to 100x improvement and highlights the potential of APT for challenging problems where PT struggles and parallel chains are constrained. By contrast, when $N\gg\Lambda$ and the improvements over standard PT are less dramatic, they might not offset the cost of additional neural evaluations. Notably, CMCD with $K=5$ and $N=30$ matches the performance of the theoretical limit $T/(2+2\Lambda)\approx 5,349$ for standard PT when $N\to\infty$ as predicted by Theorem˜2.
[…]
# Chain $N=6$ $N=10$ $N=30$
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
NF-APT 1 7.198 194 97.0 1655 827.5 2441 1220.5
CMCD-APT ($K=1$) 2 6.911 234 117.0 2126 1063.0 3264 1632.0
CMCD-APT ($K=2$) 3 5.932 526 175.3 3287 1092.7 4767 1589.0
CMCD-APT ($K=5$) 6 4.822 1743 290.5 5525 920.8 6231 1038.5
Diff-APT ($K=1$) 2 9.025 375 187.5 1551 775.5 2820 1410.0
Diff-APT ($K=2$) 3 7.298 748 249.3 2064 688.0 3480 1160.0
Diff-APT ($K=5$) 6 5.795 1565 260.8 3080 513.3 4334 722.3
Diff-PT ($K=0$) 2 8.932 204 102.0 734 367.0 1586 793.0
PT 0 8.346 17 8.5 681 340.5 1888 944.0
Figure 2: Round trip metrics for $K$-step Diff-APT ($K=1,2,5$) and Diff-PT using the true diffusion path, and Linear-PT targeting GMM-$d$ for $d=2,10,50,100$ when using 30 chains. (Left) Round trip rate against $d$. (Right) Compute-normalised round trip rate against $d$.
6.2 Scaling with Dimensions
To understand the theoretical performance of $K$-step APT for a fixed number of chains when we scale the dimension $d$, we use the fact that the path of distributions $(\pi^{\text{VP}}_{s})_{s\in[0,1]}$ induced by a VP diffusion process initialised at GMM-$d$ is analytically tractable, in order to run $K$-step Diff-APT with the true diffusion path. In Figure˜2, we compare the resulting round trip rate and compute-normalised round trip rate of the different algorithms when we fix the number of chains to 30. For all values of $d$, the round trip rate of Diff-APT monotonically increases over PT as we increase $K$, with greater gains for larger values of $d$, demonstrating that the accelerated swap improves the communication of states over the standard PT swap. We see a substantial improvement in round trip rate (right) with acceleration compared to PT ($K=0$) and a minor difference between algorithms as $K$ increases when normalised for compute (Right), suggesting the extra computation for APT is justified.
6.3 Free Energy Estimator
As we described in Section˜3.2, a by-product of Algorithm˜1 is the estimation of change in free energy $\Delta F$. We compare free energy estimates from CMCD-APT and Diff-APT against PT on two targets: DoubleWell(DW)-4, a particle system in Cartesian coordinates; and ManyWell(MW)-32, a highly multi-modal density introduced by Midgley et al. (2023). Figure˜3 presents box-plots of $30$ free energy estimates for DW-4 and MW-32 using 1,000 samples each.
Several key observations emerge: (1) Both CMCD-APT and Diff-APT exhibit markedly lower variance and bias than classical PT, across both targets. (2) For CMCD/Diff-APT, the variance and bias reduce steadily as $K$ increases. (3) For both PT and CMCD/Diff-APT, the variance and bias markedly decrease as the number of chains $N$ increases, thereby empirically validating Proposition˜1.
Figure 3: Estimates of $\Delta F$ for DW4 and ManyWell-32 by PT, CMCD-APT ($K=1,2,5$) and Diff-APT ($K=0,1,2,5$) using 1,000 samples. Each box consists of 30 estimates. The black dashed lines denotes the reference constant $\Delta F\approx 29.660$ estimated with PT using 60 chains and 100,000 samples and $\Delta F\approx 164.696$ from Midgley et al. (2023) for ManyWell-32.
6.4 Comparing APT with Neural Samplers
A significant advantage of APT is its asymptotic consistency: unlike most neural samplers, the Metropolis correction ensures APT does not incur a bias if the neural proposal is poorly trained. In Figure˜4, we demonstrate this property for both Diff-APT and CMCD-APT. Specifically, for CMCD-APT and Diff-APT, we take the trained model and map samples directly from the reference distribution to the target by concatenating the learned transports between each successive pair of intermediate distributions, labelled as CMCD and Diffusion respectively in Figure˜4.
As we can see, directly using the learned neural sampler dramatically drops performance, especially for small $K$. The diffusion model performs better than CMCD but still misallocates probability mass across two modes. On the contrary, all variants of CMCD-APT and Diff-APT recover the correct mode weights and align closely with the ground truth.
Figure 4: Interatomic distance $d_{ij}$ of 5,000 samples by CMCD, CMCD-APT, Diffusion, Diff-APT with 30 chains, $K=1,2,5$ on DW4. We take 100,000 samples by PT with 60 chains as ground truth.
Table 2: PT versus CMCD-APT targeting Alanine Dipeptide. We use the same metrics as defined in Table 1.
Method $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
CMCD-APT ($K=1$) 3.23 465 232.5
CMCD-APT ($K=2$) 3.15 597 199
CMCD-APT ($K=5$) 3.09 627 104.5
PT 3.38 199 99.5
6.5 Alanine Dipeptide
Finally, to provide a realistic target on which to evaluate APT, we take the Boltzmann distribution of Alanine Dipeptide in Cartesian coordinates at 300K, where we compare PT against CMCD-APT with $N=4$. This is a small molecule with 22 atoms (66 dimensions in total) which is highly challenging to sample from, due to many prohibitive energy barriers and high multimodality. Following the convention of molecular dynamics, we take the reference distribution to be a tempered version of the target, corresponding to 1200K. We run each algorithm for $T=50,000$ iterations and report the results in Table 2. We can see that CMCD-APT is able to provide significant acceleration in this realistic setting.
7 Conclusion
In this work, we formalise a framework integrating neural samplers to improve the sample efficiency of PT. However, several limitations, common to neural samplers, remain, including the additional burden of well optimising a neural network which has a large impact on performance - i.e. a poorly trained neural sampler may underperform PT. Further, while our approach accelerates PT by increasing the round trip rate, it incurs additional neural network evaluations. This can be computationally expensive for complex architectures. Future work should therefore develop principled criteria for deciding when to rely on PT or accelerated PT for robustness.
Acknowledgements
The authors are grateful to George Deligiannidis for helpful discussions. LZ and PP are supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1). JH acknowledges support from the University of Cambridge Harding Distinguished Postgraduate Scholars Programme. SS acknowledges support from the EPSRC CoSInES grant (EP/R034710/1) and the NSERC Postdoctoral Fellowship.
Appendix A Theoretical Analysis of Accelerated PT
A.1 Further Details of Round Trips
Consider the APT Markov chain $\mathbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ constructed by Algorithm˜1. Assume there are $m=0,\ldots,N$ parallel threads, each storing a component of $\mathbf{X}_{t}$ at time $t$. During the communication step, it is equivalent to swap indices rather than states. This has computational advantages in a distributed implementation of PT since swapping states across machines incurs a cost of $O(d)$ compared to the $O(1)$ cost of swapping indices. By tracking how the indices are shuffled, we can recover the PT state and also track the communication between reference and target through the dynamics of the permutated indices. See (Syed et al., 2022; Surjanovic et al., 2023, Algorithm 5) for details on implementing PT with a distributed implementation.
Summarising Syed et al. (2022), we define the index process $(\mathbf{I}_{t},\bm{\epsilon}_{t})$ where $\mathbf{I}_{t}=(I^{0}_{t},\ldots,I^{N}_{t})$ is a sequence of permutations of $N]=\{0,\ldots,N\}$ that tracks the underlying communication of states in $\mathbf{X}_{t}$, and $\bm{\epsilon}_{t}=(\epsilon^{0}_{t},\ldots,\epsilon^{N}_{t})$ with $\epsilon^{m}_{t}\in\{-1,1\}$ tracks the direction of the swap proposal on machine $m$.
We initialize $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ if $m$ is even and $\epsilon^{m}_{0}=1$ if $m$ is odd. The subsequent values of $\mathbf{I}_{t}$ are then determined by the swap moves. At iteration $t$ of PT, we apply the same swaps to the components of $\mathbf{I}_{t}$ that are proposed and accepted during the communication phase. As discussed in Syed et al. (2022), for non-reversible communication, $I^{m}_{t}$ and $\epsilon^{m}_{t}$ satisfy the recursion:
$I^{m}_{t}$ $=\begin{cases}I^{m}_{t-1}+\epsilon^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=1\\ I^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=0\end{cases},$
$\epsilon^{m}_{t}$ $=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$
Here $S^{n}_{t}=1$ if the $n$-th swap at time $t$ was accepted and $S^{n}_{t}=0$ otherwise.
Hence, for a realization of $\mathbf{X}_{t}$ with $T$ steps, we define the round trips for index $n\in[N]$ as the number of times the index $n$ in $\mathbf{I}_{t}$ completes the journey from the 0-th component to the $N$-th component and then back to the 0-th component—i.e., completes the round trip from the reference to the target and back again.
For $m\in[N]$, let $T^{m}_{\downarrow,0}:=\inf\{t:(I^{m}_{t},\epsilon^{m}_{t})=(0,-1)\}$ and for $j\geq 1$ define $T^{m}_{\uparrow,j}$ and $T^{m}_{\downarrow,j}$ recursively as follows:
$T^{m}_{\uparrow,j}$ $=\inf\{t>T^{m}_{\downarrow,j-1}:(I^{m}_{t},\epsilon^{m}_{t})=(N,1)\},$
$T^{m}_{\downarrow,j}$ $=\inf\{t>T^{m}_{\uparrow,j}:(I^{m}_{t},\epsilon^{m}_{t})=(0,-1)\}.$
Notably, $T^{m}_{\downarrow,j}$ represents the $j$-th time the index for machine $m$ has traversed from 0 to $N$ and back to 0, thus defining a round trip. A round trip indicates that a sample on machine $m$ from the reference propagated a reference sample to the target independent of previously visited target states on the same machine. Let $R^{m}_{T}=\max\{j:T^{m}_{\downarrow,j}\leq T\}$ denote the number of round trips that occur on machine $m$ by iteration $T$. The overall round trips count is then defined as the sum of round trips over all index values: $R_{T}=\sum_{m=0}^{N}R^{m}_{T}$. Finally, the empirical round trip rate $\tau_{T}=R_{T}/T$ represents the fraction of PT iterations during which a round trip occurred, and our objective is to maximise the expected round trip rate $\tau$ as $T\to\infty$:
$\tau:=\lim_{T\to\infty}\mathbb{E}[\tau_{T}]=\lim_{T\to\infty}\frac{\mathbb{E}[R_{T}]}{T}.$
A.2 Proof of Theorem˜1
A.2.1 Proof of Ergodicity (Theorem˜1, Part 1)
Proof.
At stationary, we have $(X_{t}^{n-1},X_{t}^{n})\sim\pi^{n-1}\otimes\pi^{n}$. We next simulate the two paths $\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1}=(\accentset{\rightharpoonup}{X}^{n-1}_{t,0},\dots,\accentset{\rightharpoonup}{X}^{n-1}_{t,K})$ and $\accentset{\leftharpoonup}{X}_{t,0:K}^{n}=(\accentset{\leftharpoonup}{X}^{n}_{t,0},\dots,\accentset{\leftharpoonup}{X}^{n}_{t,K})$ as described in Section˜3 and calculate the acceptance probability
$\alpha=\alpha^{n}_{K}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})$
given in Section˜3.2. The new states $\hat{X}_{t}^{n-1},\hat{X}_{t}^{n}$ are determined by
[…]
We would like to prove that the swap operation keeps the target distribution invariant, that is, for any real-valued test function $\varphi$ we have
$\mathbb{E}\left[\varphi(\hat{X}_{t}^{n-1},\hat{X}_{t}^{n})\right]=\mathbb{E}\left[\varphi(X_{t}^{n-1},X_{t}^{n})\right].$ (8)
One way to do this is to consider the extended target distribution
$\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1}\otimes\accentset{\leftharpoonup}{X}_{t,0:K}^{n}\sim{\mathbb{P}}_{K}^{n-1}(\mathrm{d}x_{0:K}^{n-1})\otimes{\mathbb{Q}}_{K}^{n}(\mathrm{d}x_{0:K}^{n})$
and the involution
$\mathfrak{T}(x_{0:K},y_{0:K})=(y_{0:K},x_{0:K})$
and to apply an instance of MCMC algorithms with a deterministic proposal (Tierney, 1998), here given by $\mathfrak{T}$. For completeness we give a self-contained proof. Write
$\begin{split}\mathbb{E}\left[\varphi(\hat{X}_{t}^{n-1},\hat{X}_{t}^{n})\right]&=\mathbb{E}\left[\varphi(\accentset{\leftharpoonup}{X}_{t,0}^{n},\accentset{\rightharpoonup}{X}_{t,K}^{n-1})\alpha+\varphi(\accentset{\rightharpoonup}{X}_{t,0}^{n-1},\accentset{\leftharpoonup}{X}_{t,K}^{n})(1-\alpha)\right]\\ &=\mathbb{E}\left[\varphi(X_{t}^{n-1},X_{t}^{n})\right]-\mathbb{E}\left[\varphi(\accentset{\rightharpoonup}{X}_{t,0}^{n-1},\accentset{\leftharpoonup}{X}_{t,K}^{n})\alpha\right]+\\ &\quad+\mathbb{E}\left[\varphi(\accentset{\leftharpoonup}{X}_{t,0}^{n},\accentset{\rightharpoonup}{X}_{t,K}^{n-1})\alpha\right].\end{split}$
To arrive at Equation˜8 we need to show that
$\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[\varphi\circ\psi(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\alpha\right]=\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[\varphi\circ\psi(\accentset{\leftharpoonup}{X}_{t,0:K}^{n},\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})\alpha\right]$ (9)
where $\psi(x_{0:K},y_{0:K}):=(x_{0},y_{K})$ and the notation $\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}$ means that the expectation is taken with respect to ${\mathbb{P}}_{K}^{n-1}(\mathrm{d}x_{0:K}^{n-1})\otimes{\mathbb{Q}}_{K}^{n}(\mathrm{d}x_{0:K}^{n})$. Noting that
$\alpha=\alpha^{n}_{K}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})=1\wedge\frac{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})}{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n})}$
we write
$\begin{split}&\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[\varphi\circ\psi(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\alpha\right]\\ =\ &\mathbb{E}_{{\mathbb{Q}}\otimes{\mathbb{P}}}\left[\varphi\circ\psi(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\alpha\frac{\mathrm{d}{\mathbb{P}}\otimes{\mathbb{Q}}}{\mathrm{d}{\mathbb{Q}}\otimes{\mathbb{P}}}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\right]\\ =\ &\mathbb{E}_{{\mathbb{Q}}\otimes{\mathbb{P}}}\left[\varphi\circ\psi(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\alpha\frac{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n})}{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}{(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})}}\right]\\ =\ &\mathbb{E}_{{\mathbb{Q}}\otimes{\mathbb{P}}}\left[\varphi\circ\psi(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\left\{\frac{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n})}{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})}\wedge 1\right\}\right]\\ =\ &\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[\varphi\circ\psi(\accentset{\leftharpoonup}{X}_{t,0:K}^{n},\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})\left\{\frac{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})}{\mathrm{d}{\mathbb{Q}}/\mathrm{d}{\mathbb{P}}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n})}\wedge 1\right\}\right]\\ =\ &\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[\varphi\circ\psi(\accentset{\leftharpoonup}{X}_{t,0:K}^{n},\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})\alpha\right].\end{split}$
Thus Equation˜9 is established.
Finally, the accelerated-PT Markov chain is aperiodic because the swaps can be rejected. The chain is clearly irreducible if each exploration kernel is irreducible. (In fact it can be proved, using more complicated arguments, that the chain is irreducible if the exploration kernel for the reference is irreducible and, for all $n$, the two distributions ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ are mutually absolutely continuous.) By Roberts & Rosenthal (2004, Theorem 4, Fact 5), the Markov chain is ergodic and in particular the law of large numbers hold. ∎
A.2.2 Proof of Rejection Rate (Theorem˜1, Part 2)
Proof.
The rejection rate at stationary is defined as
$r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K}):=\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}\left[1-\alpha_{K}^{n}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n})\right]$ (10)
where $\mathbb{E}_{{\mathbb{P}}\otimes{\mathbb{Q}}}$ is a shorthand for the expectation under ${\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n-1}_{K}$ and the acceptance ratio $\alpha_{K}^{n}$ is given by
$\alpha_{K}^{n}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n}):=1\wedge\frac{\mathrm{d}{\mathbb{Q}}^{n}_{K}/\mathrm{d}{\mathbb{P}}^{n-1}_{K}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1})}{\mathrm{d}{\mathbb{Q}}^{n}_{K}/\mathrm{d}{\mathbb{P}}^{n-1}_{K}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n})}=1\wedge\frac{\mathrm{d}({\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K})}{\mathrm{d}({\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K})}(\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1},\accentset{\leftharpoonup}{X}_{t,0:K}^{n}).$ (11)
From Equation˜10, Equation˜11, and Lemma˜1 below we have
$r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})=\|{\mathbb{P}}^{n-1}_{K}\otimes{\mathbb{Q}}^{n}_{K}-{\mathbb{Q}}^{n}_{K}\otimes{\mathbb{P}}^{n-1}_{K}\|_{\mathrm{TV}}.$
Lemma 1.
Let $\mu_{1}$ and $\mu_{2}$ be two mutually absolutely continuous measures. Then
$\|\mu_{1}-\mu_{2}\|_{\mathrm{TV}}=\mathbb{E}_{\mu_{1}}\left[1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right].$
Proof.
Recall the definition of the TV distance
$\|\mu_{1}-\mu_{2}\|_{\mathrm{TV}}:=\sup_{h:\mathcal{X}\to[0,1]}|\mathbb{E}_{\mu_{1}}\left[h\right]-\mathbb{E}_{\mu_{2}}\left[h\right]|.$
First, we remark that the absolute value in the definition can be omitted since
$\begin{split}\|\mu_{1}-\mu_{2}\|_{\mathrm{TV}}&=\sup_{h:\mathcal{X}\to[0,1]}\max\left(\mathbb{E}_{\mu_{1}}\left[h\right]-\mathbb{E}_{\mu_{2}}\left[h\right],\mathbb{E}_{\mu_{2}}\left[h\right]-\mathbb{E}_{\mu_{1}}\left[h\right]\right)\\ &=\sup_{h:\mathcal{X}\to[0,1]}\max\left(\mathbb{E}_{\mu_{1}}\left[h\right]-\mathbb{E}_{\mu_{2}}\left[h\right],\mathbb{E}_{\mu_{1}}\left[1-h\right]-\mathbb{E}_{\mu_{2}}\left[1-h\right]\right)\\ &=\sup_{h:\mathcal{X}\to[0,1]}\mathbb{E}_{\mu_{1}}\left[h\right]-\mathbb{E}_{\mu_{2}}\left[h\right].\end{split}$
Therefore
$\begin{split}\|\mu_{1}-\mu_{2}\|_{\mathrm{TV}}&=\sup_{h:\mathcal{X}\to[0,1]}\mathbb{E}_{\mu_{1}}\left[h\cdot\left(1-\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right]\\ &\leq\sup_{h:\mathcal{X}\to[0,1]}\mathbb{E}_{\mu_{1}}\left[h\cdot\left(1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right)\right]\\ &\leq\mathbb{E}_{\mu_{1}}\left[1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right].\end{split}$ (12)
On the other hand, let $B:=\left\{x\in X\textrm{ such that }\mathrm{d}\mu_{2}/\mathrm{d}\mu_{1}(x)\leq 1\right\}$ and put $h^{*}(x):=\mathbbm{1}_{B}(x)$. Using
$h^{*}(x)=1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}(x)\right)+\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}(x)\mathbbm{1}_{B}(x)$
write
$\begin{split}\|\mu_{1}-\mu_{2}\|_{\mathrm{TV}}&\geq\mathbb{E}_{\mu_{1}}\left[h^{*}\right]-\mathbb{E}_{\mu_{2}}\left[h^{*}\right]\\ &=\mathbb{E}_{\mu_{1}}\left[1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right]+\mathbb{E}_{\mu_{1}}\left[\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\mathbbm{1}_{B}\right]-\mathbb{E}_{\mu_{2}}\left[\mathbbm{1}_{B}\right]\\ &=\mathbb{E}_{\mu_{1}}\left[1-\min\left(1,\frac{\mathrm{d}\mu_{2}}{\mathrm{d}\mu_{1}}\right)\right].\end{split}$ (13)
Combining Equation˜12 and Equation˜13 we get the desired result. ∎
A.2.3 Proof of Proposition˜1
Proof.
We will show the consistency of the expectation and free energy estimator separately. By Theorem˜1, the PT chain $\textbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ is ergodic with stationary distribution $\pi^{0}\otimes\cdots\otimes\pi^{N}$. By the ergodic theorem, we have
$\hat{\pi}_{T}^{n}[f]=\frac{1}{T}\sum_{t=1}^{T}f(X^{n}_{t})\xrightarrow[T\to\infty]{a.s.}\pi^{n}[f].$
We will now show the consistency of the free energy estimator. By definition of $W^{n}_{K}$ in Equation˜3, we can write $\Delta F_{n}$ in terms of expectations of $W^{n}_{K}$ and the Radon-Nikodym derivative between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively:
$\exp(-\Delta F_{n})\frac{\mathrm{d}{\mathbb{Q}}^{n}_{K}}{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}=\exp(-W^{n}_{K}),\quad\exp(\Delta F_{n})\frac{\mathrm{d}{\mathbb{P}}^{n-1}_{K}}{\mathrm{d}{\mathbb{Q}}^{n}_{K}}=\exp(W^{n}_{K}).$
By taking expectations of the left and right expressions with respect to ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively, and taking the product over $n=1,\ldots,N$, we obtain the following expressions for $\exp(-\Delta F)$ and $\exp(\Delta F)$ respectively:
$\exp(-\Delta F)=\prod_{n=1}^{N}{\mathbb{P}}^{n-1}_{K}\left[\exp\left(-W^{n}_{K}\right)\right],\quad\exp(\Delta F)=\prod_{n=1}^{N}{\mathbb{Q}}^{n}_{K}\left[\exp\left(W^{n}_{K}\right)\right].$ (14)
Since $\textbf{X}_{t}$ is ergodic, $\exp(-\Delta\accentset{\rightharpoonup}{F}_{T})$ and $\exp(\Delta\accentset{\leftharpoonup}{F}_{T})$ are consistent estimators for $\exp(-\Delta F)$ and $\exp(\Delta F)$ respectively. Therefore $\Delta\hat{F}_{T}=\frac{1}{2}(\Delta\accentset{\rightharpoonup}{F}_{T}+\Delta\accentset{\leftharpoonup}{F}_{T})$ is consistent as well. Finally, note that if ${\mathbb{P}}^{n-1}_{K}={\mathbb{Q}}^{n}_{K}$, then $W^{n}_{K}(x_{0:K})\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F_{n}$, and hence $\Delta\accentset{\rightharpoonup}{F}_{T}=\Delta\accentset{\leftharpoonup}{F}_{T}\stackrel{{\scriptstyle a.s.}}{{=}}\Delta F$. ∎
A.3 Accelerated PT as Vanilla PT on Extended Space
In this section we establish the theoretical relationship between accelerated PT and vanilla PT. We state an equivalence between accelerated PT and a particular vanilla PT problem which “linearises” it. More concretely, we define a sequence of distributions $\pi_{\operatorname{ex}}^{0},\ldots,\pi_{\operatorname{ex}}^{N}$ supported on an extended space, such that if we run vanilla PT on it, we obtain the same round trip rate as if we ran accelerated PT on $\pi^{0},\ldots,\pi^{N}$.
We consider the case $K=1$ and simplify the notations of forward and backward kernels to $P^{n}(x^{n-1},\mathrm{d}x^{n})$ and $Q^{n-1}(x^{n},\mathrm{d}x^{n-1})$. This does not incur any loss of generality since conceptually multiple Markov steps can be collapsed into one.
Given a sequence of distributions $\pi^{0},\ldots,\pi^{N}$ each supported on $\mathcal{X}$, define the distributions $\pi_{\operatorname{ex}}^{n}$ as:
$\pi_{\operatorname{ex}}^{n}(\mathrm{d}x^{0},\ldots,\mathrm{d}x^{N}):=\pi^{n}(\mathrm{d}x^{n})\prod_{i\geq n+1}P^{i}(x^{i-1},\mathrm{d}x^{i})\prod_{j\leq n-1}Q^{j}(x^{j+1},\mathrm{d}x^{j}).$ (15)
In particular we stress that the distributions $\pi_{\operatorname{ex}}^{n}$ are supported on $\mathcal{X}^{N}$ and not $\mathcal{X}$.
The following proposition establishes the isometry between accelerated PT on $\pi^{1},\ldots,\pi^{N}$ and vanilla PT on $\pi_{\operatorname{ex}}^{1},\ldots,\pi_{\operatorname{ex}}^{N}$. Recall that we use $r(\mu_{1},\mu_{2})$ to denote the rejection between two distributions $\mu_{1}$ and $\mu_{2}$.
Proposition 4.
For all $1\leq n\leq N$, we have $r(\pi^{n-1}\times P^{n},Q^{n-1}\times\pi^{n})=r(\pi_{\operatorname{ex}}^{n-1},\pi_{\operatorname{ex}}^{n})$.
Proof.
Since the rejection rate only depends on the Randon-Nikodym derivative, it suffices to verify that
$\frac{\mathrm{d}\pi_{\operatorname{ex}}^{n-1}}{\mathrm{d}\pi_{\operatorname{ex}}^{n}}(x^{0},\ldots,x^{N})=\frac{\mathrm{d}(\pi^{n-1}\times P^{n})}{\mathrm{d}(Q^{n-1}\times\pi^{n})}(x^{n-1},x^{n})$
which is straightforward from Equation˜15. ∎
This proposition shows that Accelerated PT outperforms traditional PT in two ways:
  • First, while traditional PT bridges $\pi^{0}$ and $\pi^{N}$, accelerated PT bridges $\pi_{\operatorname{ex}}^{0}$ and $\pi_{\operatorname{ex}}^{N}$ which can be much closer to each other if the forward and backward kernels are good;
  • In addition, accelerated PT inserts N1N-1 distributions between $\pi_{\operatorname{ex}}^{0}$ and $\pi_{\operatorname{ex}}^{N}$.
A.4 Parallelism versus Acceleration Time
Given two distributions $\pi^{n-1}$ and $\pi^{n}$, should we apply $K$-step forward and backward kernels; or insert K1K-1 intermediate distributions $(\pi^{n-1,k})_{k=1}^{K-1}$ and use only one-step forward and backward kernels instead?
By Proposition˜2, the inverse of the local round trip rate between $\pi^{n-1}$ and $\pi^{n}$ for the first method (time-accelerated) is
$\tau^{-1}_{\textrm{TA}}:=2+2\frac{r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})}{1-r({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})}.$
The inverse of the local round trip rate between $\pi^{n-1}$ and $\pi^{n}$ for the second method (parallel-accelerated) is
$\tau^{-1}_{\textrm{PA}}:=2+2\sum_{k=1}^{K}\frac{r(\pi^{n-1,k-1}\times P^{n-1}_{k},Q^{n}_{k-1}\times\pi^{n-1,k})}{1-r(\pi^{n-1,k-1}\times P^{n-1}_{k},Q^{n}_{k-1}\times\pi^{n-1,k})}$
where we make the convention $\pi^{n-1,0}\equiv\pi^{n-1}$ and $\pi^{n-1,K}\equiv\pi^{n}$.
The following proposition analyses these local rates for both methods. As in Section˜A.3, the main idea is to find a vanilla PT equivalent for both algorithms. We define
$\tau^{-1}_{\textrm{VA}}(\mu_{0},\ldots,\mu_{K}):=2+\sum_{k=1}^{K}\frac{r(\mu_{k-1},\mu_{k})}{1-r(\mu_{k-1},\mu_{k})}$
as the inverse round trip rate of a vanilla PT algorithm on a sequence of distributions $\mu_{0},\ldots,\mu_{K}$.
Proposition 5.
Define the sequence of distributions $({\mathbb{S}}_{k})_{k=0}^{K}$ as
${\mathbb{S}}_{k}(\mathrm{d}x_{0},\mathrm{d}x_{1},\ldots,\mathrm{d}x_{K}):=\pi^{n-1,k}(\mathrm{d}x_{k})\times\prod_{i\geq k+1}P^{n-1}_{i}(x_{i-1},\mathrm{d}x_{i})\prod_{j\leq k-1}Q^{n}_{j}(x_{j+1},\mathrm{d}x_{j}).$
Then the following equalities hold
$\tau_{\textrm{TA}}$ $=\tau_{\textrm{VA}}({\mathbb{S}}_{0},{\mathbb{S}}_{K})$ (16)
$\tau_{\textrm{PA}}$ $=\tau_{\textrm{VA}}({\mathbb{S}}_{0},{\mathbb{S}}_{1},\ldots,{\mathbb{S}}_{K}).$ (17)
Proof.
The first point is straightforward. To show the second point, we need to check that
$r(\pi^{n-1,k-1}\times P^{n-1}_{k},Q^{n}_{k-1}\times\pi^{n-1,k})=r({\mathbb{S}}_{k-1},{\mathbb{S}}_{k}).$
Note that the rejection rates only depend on the Radon-Nikodym derivatives, so it suffices to verify that
$\frac{\mathrm{d}{\mathbb{S}}_{k}}{\mathrm{d}{\mathbb{S}}_{k-1}}(x_{0},x_{1},\ldots,x_{K})=\frac{\mathrm{d}(Q^{n}_{k-1}\times\pi^{n-1,k})}{\mathrm{d}(\pi^{n-1,k-1}\times P^{n-1}_{k})}(x_{k-1},x_{k})$
which is straightforward from the definition of $({\mathbb{S}}_{k})_{k=0}^{K}$. ∎
This proposition shows that it is preferable to use the parallel-accelerated method, as the quantity in Equation˜17 is generally greater than that of Equation˜16 thanks to the effect of the bridge between ${\mathbb{S}}_{0}$ and ${\mathbb{S}}_{K}$. However, in practice the time-accelerated method consumes less memory and so might be more suitable in certain circumstances.
Appendix B Further Details on Analysis of PT
B.1 Proof of Proposition˜2
Let $S^{n}_{t}$ be the indicator random variable with $S^{n}_{t}=1$ if the $n$-th swap is proposed and accepted at iteration $t$, and $S^{n}_{t}=0$ otherwise. By ˜1, the random variables $S^{n}_{t}$ are i.i.d. in $t$ with $S^{n}_{1},S^{n}_{2},\ldots\stackrel{{\scriptstyle d}}{{=}}\mathrm{Bernoulli}(s_{n})$, where $s_{n}:=1-r({\mathbb{P}}^{n-1},{\mathbb{Q}}^{n})$. This implies that the index processes $(I^{m}_{t},\epsilon^{m}_{t})$ form Markov chains on $\{0,\ldots,N\}\times\{-1,1\}$ with initial conditions $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ when $m$ is even, $\epsilon^{m}_{0}=1$ when $m$ is odd. For $t\geq 1$, each process satisfies the following recursion:
$I^{m}_{t}$ $\sim\begin{cases}I^{m}_{t-1}+\epsilon^{m}_{t-1},&\text{with probability }s_{I^{m}_{t-1}\vee(I^{m}_{t-1}+\epsilon^{m}_{t-1})}\\ I^{m}_{t-1},&\text{with probability }1-s_{I^{m}_{t-1}\vee(I^{m}_{t-1}+\epsilon^{m}_{t-1})}\end{cases},$ (18)
$\epsilon^{m}_{t}$ $=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$ (19)
The remainder of the proof follows identically to (Syed et al., 2022, Corollary 2).
B.2 Proof of Proposition˜3
We first spell out the full statement of Proposition˜3.
Proposition.
Suppose that ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ arise as $K$-step Euler-Maruyama discretizations of the following SDEs respectively
$d\accentset{\rightharpoonup}{X}^{n-1}_{t}$ $=f(t,\accentset{\rightharpoonup}{X}^{n-1}_{t})dt+\sigma dB_{t},$ (20)
$d\accentset{\leftharpoonup}{X}^{n}_{t}$ $=b(t,\accentset{\leftharpoonup}{X}^{n}_{t})dt+\sigma dB_{t},$ (21)
[…]
$\lim_{K\to\infty}r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})$ (22)
and
$r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})\leq r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})+\mathcal{O}(\frac{1}{\sqrt{K}})$ (23)
where $r(p,q)=\|p\times q-q\times p\|_{\mathrm{TV}}$ and ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ are respectively the full path measures of Equation˜20 and Equation˜21.
Proof.
Let ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ be the restrictions of the path measures ${\mathbb{P}}_{\infty}^{n-1}$ and ${\mathbb{Q}}_{\infty}^{n}$ to the time discretization points. (As such, ${\mathbb{P}}_{\infty|K}^{n-1}$ and ${\mathbb{Q}}_{\infty|K}^{n}$ are defined on the same space as ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$.) We have
$\begin{split}\quad|r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})-r({\mathbb{P}}_{\infty|K}^{n-1},{\mathbb{Q}}_{\infty|K}^{n})|&\leq 2\left(\|{\mathbb{P}}_{K}^{n-1}-{\mathbb{P}}_{\infty|K}^{n-1}\|_{\mathrm{TV}}+\|{\mathbb{Q}}_{K}^{n}-{\mathbb{Q}}_{\infty|K}^{n}\|_{\mathrm{TV}}\right)\\ &\leq\mathcal{O}(1/\sqrt{K})\end{split}$ (24)
where the first inequality is elementary and the second follows from Proposition˜6. By data processing inequality
$r({\mathbb{P}}_{\infty|K}^{n-1},{\mathbb{Q}}_{\infty|K}^{n})\leq r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n}).$ (25)
Equation˜24 and Equation˜25 establish Equation˜23. On the other hand, putting
$R_{K}(x_{k})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{K}^{n}\times{\mathbb{P}}_{K}^{n-1})}{\mathrm{d}({\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n})}(x_{k});$
$R_{\infty}(x_{\infty})$ $:=\frac{\mathrm{d}({\mathbb{Q}}_{\infty}^{n}\times{\mathbb{P}}_{\infty}^{n-1})}{\mathrm{d}({\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n})}(x_{\infty});$
$\varphi(\alpha)$ $:=1-\min(1,\alpha)$
we have by Lemma˜1
$r({\mathbb{P}}_{K}^{n-1},{\mathbb{Q}}_{K}^{n})=\mathbb{E}_{{\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n}}[\varphi\circ R_{K}].$ (26)
Since $\varphi(\alpha)\in[0,1]$ the definition of the total variation distance implies
$\left|\mathbb{E}_{{\mathbb{P}}_{K}^{n-1}\times{\mathbb{Q}}_{K}^{n}}[\varphi\circ R_{K}]-\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]\right|\leq\mathcal{O}(1/\sqrt{K}).$ (27)
In addition let $D_{K}:\mathcal{C}[0,1]\to(\mathbb{R}^{d})^{K+1}$ be the discretisation operator which takes a continuous path $x_{\infty}$ and extract its values at $K+1$ discretisation points. Then
$\begin{split}\mathbb{E}_{{\mathbb{P}}_{\infty|K}^{n-1}\times{\mathbb{Q}}_{\infty|K}^{n}}[\varphi\circ R_{K}]&=\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{K}\circ D_{K}]\\ &\overset{K\to\infty}{\longrightarrow}\mathbb{E}_{{\mathbb{P}}_{\infty}^{n-1}\times{\mathbb{Q}}_{\infty}^{n}}[\varphi\circ R_{\infty}]=r({\mathbb{P}}_{\infty}^{n-1},{\mathbb{Q}}_{\infty}^{n})\end{split}$ (28)
by dominated convergence theorem, taking into account the pathwise convergence of the Euler-Maruyama schemes (Kloeden & Neuenkirch (2007); Gyöngy (1998), see also Berner et al. (2025, Lemma B.7.a)). Equation˜26, Equation˜27, and Equation˜28 entail Equation˜22, finishing the proof.
Proposition 6.
Let $(X_{t})_{t\in[0,1]}$ be the solution of an SDE of the form
$dX_{t}=b(t,X_{t})dt+\sigma dB_{t},$ (29)
where we assume that $b(t,\cdot)$ is $L$-Lipschitz continuous for all $t$ and that $b(\cdot,x)$ is $B$-Lipschitz continuous for all $x$. We also assume a linear growth condition of the form $\|b(t,x)\|\leq G(1+\|x\|)$. Consider the $K$-step Euler-Maruyama discretization $\hat{X}_{t_{0}},\hat{X}_{t_{1}},\dots,\hat{X}_{t_{K}}$ where $t_{k}=k/K$. Let $\mathbb{P}$$_{\infty}$ be the path-measure of the process $(X_{t})_{t\in[0,1]}$, ${\mathbb{P}}^{*}_{K}$ be the law of ${X}_{t_{0}},\dots,{X}_{t_{K}}$, and ${{\mathbb{P}}}_{K}$ be the law of $\hat{X}_{t_{0}},\dots,\hat{X}_{t_{K}}$. Then
$\|{\mathbb{P}}^{*}_{K}-{{\mathbb{P}}}_{K}\|_{\mathrm{TV}}\in\mathcal{O}(\frac{1}{\sqrt{K}}).$
Proof.
Consider the continuous-time extension $(\hat{X}_{t})_{t\in[0,1]}$ of the $K$-step Euler-Maruyama discretization of the SDE in Equation˜29 given by
$d\hat{X}_{t}=b(t_{k},\hat{X}_{t_{k}})dt+\sigma dB_{t},\quad t\in[t_{k},t_{k+1}).$
We use $\hat{{\mathbb{P}}}_{\infty}=\hat{{\mathbb{P}}}^{(K)}_{\infty}$ to denote its path-measure. By Pinsker’s inequality
$\|{\mathbb{P}}^{*}_{K}-{{\mathbb{P}}}_{K}\|_{\mathrm{TV}}$ $\leq\sqrt{\frac{1}{2}\text{KL}({\mathbb{P}}^{*}_{K}\|{{\mathbb{P}}}_{K})}.$
By the data-processing inequality and Girsanov’s theorem, we have that
$\text{KL}({\mathbb{P}}^{*}_{K}\|{{\mathbb{P}}}_{K})\leq\text{KL}({{\mathbb{P}}}_{\infty}\|\hat{{\mathbb{P}}}_{\infty})=\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}\|b(t,X_{t})-b(t_{k},X_{t_{k}})\|^{2}dt.$
By the Lipschitz continutity of $b$, we have the following bound
$\mathbb{E}\|b(t,X_{t})-b(t_{k},X_{t_{k}})\|^{2}$ $\leq 2\mathbb{E}\|b(t,X_{t})-b(t,X_{t_{k}})\|^{2}+2\mathbb{E}\|b(t,X_{t_{k}})-b(t_{k},X_{t_{k}})\|^{2}$
$\leq 2L^{2}\mathbb{E}\|X_{t}-X_{t_{k}}\|^{2}+2B^{2}(t-t_{k})^{2}\mathbb{E}\|X_{t_{k}}\|^{2}.$
By the Cauchy-Schwarz inequality, the fact that $B_{t}-B_{t_{k}}\sim\mathcal{N}(0,(t-t_{k})I_{d})$ and the linear growth assumption, we obtain
$\mathbb{E}\|X_{t}-X_{t_{k}}\|^{2}$ $=\mathbb{E}\left\|\int_{t_{k}}^{t}b(s,X_{s})ds+\sigma(B_{t}-B_{t_{k}})\right\|^{2}$
$\leq 2(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}\left\|b(s,X_{s})\right\|^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 2G(t-t_{k})\int_{t_{k}}^{t}\mathbb{E}(1+\|X_{s}\|)^{2}ds+2d\sigma^{2}(t-t_{k})$
$\leq 4G(t-t_{k})\int_{t_{k}}^{t}(1+\mathbb{E}\|X_{s}\|^{2})ds+2d\sigma^{2}(t-t_{k})$
It is well known that under the assumptions in the proposition, there exists a constant $M$ (independent of $K$) such that $\mathbb{E}\|X_{t}\|^{2}\leq M$ for all $t$. Putting all of this together, we have shown that
$\mathbb{E}\|b(t,X_{t})-b(t_{k},X_{t_{k}})\|^{2}$ $\leq 8GL^{2}(1+M)(t-t_{k})^{2}+4d\sigma^{2}L^{2}(t-t_{k})+2MB^{2}(t-t_{k})^{2}.$
Summing over $k$ and integrating, this yields
$\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}\|b(t,X_{t})-b(t_{k},X_{t_{k}})\|^{2}dt$
$\leq\frac{1}{2\sigma^{2}}\sum_{k=0}^{K-1}\frac{8}{3}GL^{2}(1+M)(t_{k+1}-t_{k})^{3}+2d\sigma^{2}L^{2}(t_{k+1}-t_{k})^{2}+\frac{2}{3}MB^{2}(t_{k+1}-t_{k})^{3}$
$\leq\frac{4}{3\sigma^{2}}GL^{2}(1+M)\frac{1}{K^{2}}+dL^{2}\frac{1}{K}+\frac{1}{3\sigma^{2}}MB^{2}\frac{1}{K^{2}}.$
Therefore
$\|{\mathbb{P}}^{*}_{K}-{{\mathbb{P}}}_{K}\|_{\mathrm{TV}}\in\mathcal{O}(1/\sqrt{K}).$
We verify this result in practice by plotting the TV distance between ${\mathbb{P}}_{K}$ and $\mathbb{P}$$_{\infty}$ for the Ornstein–Uhlenbeck process in 1 dimension.
B.3 Further Details on Scaling with Parallel Chains
B.3.1 Regularity Assumptions
Suppose $\pi^{\beta}(x)=\exp(-U^{\beta}(x))/Z_{\beta}$, where $Z_{\beta}=\int_{\mathcal{X}}\exp(-U^{\beta}(x))\mathrm{d}x$. For $\bm{\beta}=(\beta,\beta^{\prime})$, suppose ${\mathbb{P}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ and ${\mathbb{Q}}^{\bm{\beta}}_{K}(\mathrm{d}x_{0:K})$ are mutually absolutely continuous path measures on $\mathcal{X}^{K+1}$ with marginals ${\mathbb{P}}^{\bm{\beta}}(x_{0})=\pi^{\beta}(\mathrm{d}x_{0})$, and ${\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{K})=\pi^{\beta^{\prime}}(\mathrm{d}x_{K})$ with work functional $W^{\bm{\beta}}:\mathcal{X}^{K+1}\mapsto\mathbb{R}$,
$W^{\bm{\beta}}(x_{0:K})=\Delta F_{\bm{\beta}}-\log\frac{\mathrm{d}{\mathbb{P}}^{\bm{\beta}}}{\mathrm{d}{\mathbb{Q}}^{\bm{\beta}}}(x_{0:K}),$
with $\Delta F_{\bm{\beta}}:=\log Z_{\beta^{\prime}}-\log Z_{\beta}$.
We will make the following regularity assumptions on ${\mathbb{P}}^{\bm{\beta}}$ and ${\mathbb{Q}}^{\bm{\beta}}$ analogous to Assumptions 5–7 in Syed et al. (2024). Suppose there exists $\mathcal{F}=(\mathcal{F}_{i})_{i\in\mathbb{N}}$ of a nested sequence $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\cdots$ of vector spaces of measurable function $f:\mathcal{X}^{K+1}\to\mathbb{R}$, such that (1) $\mathcal{F}_{i}$ contains the constant functions, and (2) $\mathcal{F}_{i}$ is closed under domination, i.e., if $g$ is a measurable function such that $|g|\leq|f|$ for some $f\in\mathcal{F}_{i}$, then $g\in\mathcal{F}_{i}$.
Assumption 2.
For $i=0,1,2$, and $\bm{\beta}\mapsto{\mathbb{P}}^{\bm{\beta}}[f]$ and $\bm{\beta}\mapsto{\mathbb{Q}}^{\bm{\beta}}[f]$ are $i$-times continuously differentiable for all $f\in\mathcal{F}_{i}$, there exists signed-measures $\partial_{\bm{\beta}}^{i}{\mathbb{P}}^{\bm{\beta}}$ and $\partial_{\bm{\beta}}^{i}{\mathbb{Q}}^{\bm{\beta}}$ integrable over $\mathcal{F}_{i}$ such that for all $f\in\mathcal{F}_{i}$,
$\partial^{i}_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}[f]=\partial_{\bm{\beta}}^{i}({\mathbb{P}}^{\bm{\beta}}[f]),\quad\partial^{i}_{\bm{\beta}}{\mathbb{Q}}^{\bm{\beta}}[f]=\partial_{\bm{\beta}}^{i}({\mathbb{Q}}^{\bm{\beta}}[f]).$
Assumption 3.
$\bm{\beta}\mapsto W^{\bm{\beta}}$ is $2$-times continuously differentiable, and for $i=0,1,2$ and $i_{1},\dots,i_{j}\geq 1$ and $i_{1}+\dots+i_{j}=i$, we have $\bar{W}_{i_{1}}\cdots\bar{W}_{i_{j}}\in\mathcal{F}_{3-i}$, where $\bar{W}_{i}=\sup_{\bm{\beta}}\left|\partial_{\bm{\beta}}^{i}W^{\bm{\beta}}\right|$.
Assumption 4.
For all $\beta$ we have ${\mathbb{P}}^{\beta}:={\mathbb{P}}^{\beta,\beta}={\mathbb{Q}}^{\beta,\beta}=:{\mathbb{Q}}^{\beta}$.
Lemma 2.
Let $g^{\bm{\beta}}:\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}\to\mathbb{R}$,
  • $\bm{\beta}\mapsto g^{\bm{\beta}}(x_{0:K},x_{0:K})$ is continuously differentiable,
  • $x_{0:K}\mapsto g^{\bm{\beta}}(x_{0:K},x^{\prime}_{0:K})$ and $x_{0:K}^{\prime}\mapsto g^{\bm{\beta}}(x_{0:K},x^{\prime}_{0:K})$ are in $\mathcal{F}_{1}$,
  • $x_{0:K}\mapsto\partial_{\bm{\beta}}g^{\bm{\beta}}(x_{0:K},x^{\prime}_{0:K})$ and $x_{0:K}^{\prime}\mapsto\partial_{\bm{\beta}}g^{\bm{\beta}}(x_{0:K},x^{\prime}_{0:K})$ are in $\mathcal{F}_{0}$.
Then $\bm{\beta}\mapsto{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}]$ is continuously differentiable with derivative,
$\partial_{\bm{\beta}}({\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}])=\partial_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[g]+{\mathbb{P}}^{\bm{\beta}}\otimes\partial_{\bm{\beta}}{\mathbb{Q}}^{\bm{\beta}}[g]+{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\partial_{\bm{\beta}}g^{\bm{\beta}}].$
Proof.
For notational convenience we denote $x:=x_{0:K}$ and $x^{\prime}:=x_{0:K}$. Given $x\in\mathcal{X}^{K+1}$ let $g_{x}^{\bm{\beta}}:\mathcal{X}^{K+1}\to\mathbb{R}$ be the marginal $g^{\bm{\beta}}_{x}(x^{\prime})=g(x,x^{\prime})$ and let $q^{\bm{\beta}}:\mathcal{X}^{K+1}\to\mathbb{R}$ denote the expectation of $g_{x}^{\bm{\beta}}$ with respect to ${\mathbb{Q}}^{\bm{\beta}}$, i.e. $q^{\bm{\beta}}(x)={\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}_{x}]$. By product rule for measure-valued derivatives, we have for all $x$, $\bm{\beta}\mapsto q^{\bm{\beta}}(x)$ is continuously differentiable with derivative,
$\partial_{\bm{\beta}}q^{\bm{\beta}}(x)$ $=\partial_{\bm{\beta}}({\mathbb{Q}}^{\bm{\beta}}[g_{x}^{\bm{\beta}}])$
$=\partial_{\bm{\beta}}{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}_{x}]+{\mathbb{Q}}^{\bm{\beta}}[\partial_{\bm{\beta}}g^{\bm{\beta}}_{x}].$
by taking expectation of both sides with respect to ${\mathbb{P}}^{\bm{\beta}}$ and using Fubini’s theorem we have
${\mathbb{P}}^{\bm{\beta}}[\partial_{\bm{\beta}}q^{\bm{\beta}}]={\mathbb{P}}^{\bm{\beta}}\otimes\partial_{\bm{\beta}}{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}]+{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\partial_{\bm{\beta}}g^{\bm{\beta}}].$
Again, using the product rule for measure-valued derivatives,
$\partial_{\bm{\beta}}({\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}])$ $=\partial_{\bm{\beta}}({\mathbb{P}}^{\bm{\beta}}[q^{\bm{\beta}}])$
$=\partial_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}[q^{\bm{\beta}}]+{\mathbb{P}}^{\bm{\beta}}[\partial_{\bm{\beta}}q^{\bm{\beta}}].$
The result follows by noting that $\partial_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}[q^{\bm{\beta}}]=\partial_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[g^{\bm{\beta}}]$ by Fubini’s theorem. ∎
B.3.2 Proof of Theorem˜2
For $\bm{\beta}=(\beta,\beta^{\prime})$ recall the rejection rate can be expressed as $r({\mathbb{P}}^{\bm{\beta}},{\mathbb{Q}}^{\bm{\beta}})=1-{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\alpha^{\bm{\beta}}]$, where ${\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}$ is a measure over $\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}$ is the product between the forward and backwards paths,
${\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{0:K},\mathrm{d}x_{0:K}^{\prime}):={\mathbb{P}}^{\bm{\beta}}(\mathrm{d}x_{0:K}){\mathbb{Q}}^{\bm{\beta}}(\mathrm{d}x_{0:K}^{\prime}),$
$\alpha^{\bm{\beta}}:\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}\to[0,1]$ is the acceptance probability for swap proposed,
$\alpha^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime})$ $=\exp\left(\min\left\{0,\Delta W^{\bm{\beta}}_{K}(x_{0:K},x_{0:K}^{\prime})\right\}\right)$
and $\Delta W^{\bm{\beta}}:\mathcal{X}^{K+1}\times\mathcal{X}^{K+1}\to\mathbb{R}$ is the change in work,
$\Delta W^{\bm{\beta}}(x_{0:K},x_{0:K}^{\prime}):=W^{\bm{\beta}}(x_{0:K}^{\prime})-W^{\bm{\beta}}(x_{0:K}).$
Lemma 3.
Suppose Assumptions 2–4 hold. For all $\bm{\beta}=(\beta,\beta^{\prime})$ with $\Delta\beta=\beta^{\prime}-\beta>0$, exists a constant $C>0$ independent of $\bm{\beta}$ such that,
$\left|r({\mathbb{P}}^{\bm{\beta}},{\mathbb{Q}}^{\bm{\beta}})-\int_{\beta}^{\beta^{\prime}}\lambda_{b}\mathrm{d}b\right|\leq C\Delta\beta^{2}.$
where $\lambda_{\beta}:=\frac{1}{2}{\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[|\Delta\dot{W}^{\beta}|]$ and $\Delta\dot{W}^{\beta}:=\lim_{\beta^{\prime}\to\beta}\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}$. Moreover, $\lambda_{\beta}$ is continuously differentiable in $\beta$.
Given an annealing schedule $0=\beta_{0}<\cdots<\beta_{N}=1$, let $\bm{\beta}_{n}=(\beta_{n-1},\beta_{n})$ and $\Delta\beta_{n}=\beta_{n}-\beta_{n-1}$. It follows from Lemma˜3 that the sum of the rejection satisfies,
$\left|\sum_{n=1}^{N}r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})-\Lambda\right|\leq C\sum_{n=1}^{N}\Delta\beta_{n}^{2}\leq C\max_{n\leq N}|\Delta\beta_{n}|.$
Therefore as $N\to\infty$ and $\max_{n\leq N}|\Delta\beta_{n}|\to 0$ we have the sum of the rejection rates converge to $\Lambda$.
For the same annealing schedule define $\tau^{N}(\beta_{0:N})$ as
$\tau^{N}(\beta_{0:N}):=\left(2+2\sum_{n=1}^{N}\frac{r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}{1-r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}\right)^{-1}$
By Lemma˜3 we have,
$r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})\leq\frac{r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}{1-r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}\leq\frac{r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}{1-\max_{n\leq N}r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}\leq\frac{r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}{1-\sup_{\beta}\lambda_{\beta}\max_{n\leq N}|\Delta\beta_{n}|}.$
Therefore, by taking the sum over $n$ and using the squeeze theorem, we have in the limit $N\to\infty$ and $\max_{n\leq N}|\Delta\beta_{n}|\to 0$,
$\lim_{N\to\infty}\sum_{n=1}^{N}\frac{r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}{1-r({\mathbb{P}}^{\bm{\beta}_{n}},{\mathbb{Q}}^{\bm{\beta}_{n}})}=\Lambda$
and hence $\lim_{N\to\infty}\tau^{N}(\beta_{0:N})=(2+2\Lambda)^{-1}$, which completes the proof.
Proof of Lemma˜3.
Since $\bm{\beta}\mapsto W^{\bm{\beta}}$ is twice differentiable, we have that $\bm{\beta}\mapsto\alpha^{\bm{\beta}}$ is twice differentiable when $\Delta W^{\bm{\beta}}\neq 0$, with first-order partial derivatives with respect to $\beta^{\prime}$:
$\partial_{\beta^{\prime}}\alpha^{\bm{\beta}}$ $=\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}\exp(\Delta W^{\bm{\beta}})1[\Delta W^{\bm{\beta}}<0],$
and the second-order partial derivative with respect to $\beta^{\prime}$:
$\partial_{\beta^{\prime}}^{2}\alpha^{\bm{\beta}}$ $=[\partial_{\beta^{\prime}}^{2}\Delta W^{\bm{\beta}}+(\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}})^{2}]\exp(\Delta W^{\bm{\beta}})1[\Delta W^{\bm{\beta}}<0].$
By Taylor’s theorem, for $\Delta\beta>0$, we have
$\alpha^{\bm{\beta}}=1+\dot{\alpha}^{\beta}\Delta\beta+\epsilon^{\bm{\beta}},$
where $\dot{\alpha}^{\beta}=\lim_{\beta^{\prime}\to\beta^{+}}\partial_{\beta^{\prime}}\alpha^{\bm{\beta}}$ and $|\epsilon^{\bm{\beta}}|\leq\frac{1}{2}\sup_{\bm{\beta}}|\partial_{\beta^{\prime}}^{2}\alpha^{\bm{\beta}}|\Delta\beta^{2}$. Therefore, the rejection rate equals:
$r({\mathbb{P}}^{\bm{\beta}},{\mathbb{Q}}^{\bm{\beta}})=-{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]\Delta\beta-{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\epsilon^{\bm{\beta}}].$ (30)
We will first approximate the first term in Equation˜30. Note that since $\Delta W^{\beta}=0$ and $\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}=\lim_{\beta^{\prime}\to\beta}\Delta W^{\bm{\beta}}/\Delta\beta=:\Delta\dot{W}^{\bm{\beta}}$, we have that $\dot{\alpha}^{\beta}$ satisfies:
$\dot{\alpha}^{\beta}$ $=\lim_{\beta^{\prime}\to\beta}\partial_{\beta^{\prime}}\Delta W^{\bm{\beta}}\exp(\Delta W^{\bm{\beta}})1\left[\frac{\Delta W^{\bm{\beta}}}{\Delta\beta}<0\right]$
$=\Delta\dot{W}^{\beta}1[\Delta\dot{W}^{\beta}<0]$
$=-|\Delta\dot{W}^{\beta}|1[\Delta\dot{W}^{\beta}<0].$
We can bound $\dot{\alpha}^{\beta}$ uniformly in $\beta$ in terms of $\bar{W}_{1}$:
$|\dot{\alpha}^{\beta}(x_{0:K},x_{0:K}^{\prime})|$ $\leq|\dot{W}^{\beta}(x_{0:K})-\dot{W}^{\beta}(x_{0:K}^{\prime})|$
$\leq\bar{W}_{1}(x_{0:K})+\bar{W}_{1}(x_{0:K}^{\prime})$
$:=\bar{W}_{1}\oplus\bar{W}_{1}(x_{0:K},x_{0:K}^{\prime}).$
[…]
$\frac{{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]-{\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[\dot{\alpha}^{\beta}]}{\Delta\beta}$ $=\partial_{\beta^{\prime}}({\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}])|_{\bm{\beta}=\tilde{\bm{\beta}}}$
$=\partial_{\beta^{\prime}}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]|_{\bm{\beta}=\tilde{\bm{\beta}}}+{\mathbb{P}}^{\bm{\beta}}\otimes\partial_{\beta^{\prime}}{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]|_{\bm{\beta}=\tilde{\bm{\beta}}}$.
Using the triangle inequality and $|\dot{\alpha}^{\beta}|\leq\bar{W}_{1}\oplus\bar{W}_{1}$, we have:
$\left|\frac{{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]-{\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[\dot{\alpha}^{\beta}]}{\Delta\beta}\right|$ $\leq\sup_{\bm{\beta}}|\partial_{\beta^{\prime}}{\mathbb{P}}^{\bm{\beta}}|\otimes{\mathbb{Q}}^{\bm{\beta}}[\bar{W}_{1}\oplus\bar{W}_{1}]$
$\quad+\sup_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}\otimes|\partial_{\beta^{\prime}}{\mathbb{Q}}^{\bm{\beta}}|[\bar{W}_{1}\oplus\bar{W}_{1}]$
$=\sup_{\bm{\beta}}\left(|\partial_{\beta^{\prime}}{\mathbb{P}}^{\bm{\beta}}|[1]{\mathbb{Q}}^{\bm{\beta}}[\bar{W}_{1}]+|\partial_{\beta^{\prime}}{\mathbb{P}}^{\bm{\beta}}|[\bar{W}_{1}]{\mathbb{Q}}^{\bm{\beta}}[1]\right)$
$\quad+\sup_{\bm{\beta}}\left({\mathbb{P}}^{\bm{\beta}}[1]|\partial_{\beta^{\prime}}{\mathbb{Q}}^{\bm{\beta}}|[\bar{W}_{1}]+{\mathbb{P}}^{\bm{\beta}}[\bar{W}_{1}]|\partial_{\beta^{\prime}}{\mathbb{Q}}^{\bm{\beta}}|[1]\right).$
Since $1$ and $\bar{W}_{1}$ are in $\mathcal{F}_{1}$, we have that each of the terms on the right-hand side is continuous and hence by the extreme value theorem, there exists $C_{1}$ such that
$|{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\dot{\alpha}^{\beta}]-{\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[\dot{\alpha}^{\beta}]|\leq C_{1}\Delta\beta.$
For the second term in Equation˜30, we have
$|{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\epsilon^{\bm{\beta}}]|$ $\leq\frac{\Delta\beta^{2}}{2}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\sup_{\bm{\beta}}|\partial_{\beta^{\prime}}^{2}\alpha^{\bm{\beta}}|]$
$\leq\frac{\Delta\beta^{2}}{2}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\bar{W}_{2}\oplus\bar{W}_{2}+(\bar{W}_{1}\oplus\bar{W}_{1})^{2}]$
$\leq\frac{\Delta\beta^{2}}{2}\sup_{\bm{\beta}}{\mathbb{P}}^{\bm{\beta}}\otimes{\mathbb{Q}}^{\bm{\beta}}[\bar{W}_{2}\oplus\bar{W}_{2}+(\bar{W}_{1}\oplus\bar{W}_{1})^{2}]$
$:=C_{2}\Delta\beta^{2},$
where $\bar{W}_{2}\oplus\bar{W}_{2}(x_{0:K},x_{0:K}^{\prime}):=\bar{W}_{2}(x_{0:K})+\bar{W}_{2}(x_{0:K}^{\prime})$. ˜2 guarantees that the expectation in the second-to-last line is continuous in $\bm{\beta}$, and hence $C_{2}$ is finite. Next, we note that since $|\Delta\dot{W}^{\beta}(x_{0:K},x_{0:K}^{\prime})|=|\Delta\dot{W}^{\beta}(x_{0:K}^{\prime},x_{0:K})|$ is symmetric and ${\mathbb{P}}^{\beta}={\mathbb{Q}}^{\beta}$, we have:
${\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[\dot{\alpha}^{\beta}]$ $={\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[-|\Delta\dot{W}^{\beta}|1[\Delta\dot{W}^{\beta}<0]]$
$=-\frac{1}{2}{\mathbb{P}}^{\beta}\otimes{\mathbb{Q}}^{\beta}[|\Delta\dot{W}^{\beta}|]$
$=-\lambda_{\beta}.$
By Lemma˜2, $\beta\mapsto\lambda_{\beta}$ is continuously differentiable. Since $\lambda_{\beta}\Delta\beta$ is a right Riemann sum for the integral of $\lambda_{\beta}$ with error:
$\left|\lambda_{\beta}\Delta\beta-\int_{\beta}^{\beta^{\prime}}\lambda_{b}\,\mathrm{d}b\right|=\frac{1}{2}\sup_{\beta^{\prime}}\left|\frac{\mathrm{d}\lambda_{\beta}}{\mathrm{d}\beta}\right|\Delta\beta^{2}=:C_{3}\Delta\beta^{2}.$
Finally, by the triangle inequality:
$\left|r({\mathbb{P}}^{\bm{\beta}},{\mathbb{Q}}^{\bm{\beta}})-\int_{\beta}^{\beta^{\prime}}\lambda_{b}\,\mathrm{d}b\right|\leq C\Delta\beta^{2},$
for $C:=C_{1}+C_{2}+C_{3}$. ∎
Appendix C Further Details on Design-Space for Accelerated PT
C.1 Further Details on Flow APT
C.1.1 Work Formula for Deterministic flows
Proposition 7.
Let $\mathcal{X}=\mathbb{R}^{d}$ and suppose that $\pi^{n-1}$ and $\pi^{n}$ admit strictly positive densities $\tilde{\pi}^{n-1}$ and $\tilde{\pi}_{n}$ with respect to the Lebesgue measure. Let $T^{n}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a diffeomorphism with Jacobian matrix $J_{T_{n}}(x)$. If we choose the one-step forward and backward kernels $P^{n-1}$ and $Q^{n}$ such that
$P^{n-1}(x^{n-1},\mathrm{d}x^{n}_{*})=\delta_{T^{n}(x^{n-1})}(\mathrm{d}x^{n}_{*}),\qquad Q^{n}(x^{n},\mathrm{d}x^{n-1}_{*})=\delta_{(T^{n})^{-1}(x^{n})}(\mathrm{d}x^{n-1}_{*}),$
then, for all $(z^{n-1},z^{n})$ such that $T^{n}(z^{n-1})=z^{n}$, we have the following expression for the work defined in Equation˜3:
$W^{n}(z^{n-1},z^{n})=U^{n}(z^{n})-U^{n-1}(z^{n-1})-\log|\det J_{T^{n}}(z^{n-1})|;$
and the acceptance rate $\alpha^{n}$ defined in Section˜3.2 becomes
$\alpha^{n}(x^{n-1},x^{n};x^{n-1}_{*},x^{n}_{*})=1\wedge\left[\frac{\tilde{\pi}^{n-1}(x^{n-1}_{*})\tilde{\pi}^{n}(x^{n}_{*})}{\tilde{\pi}^{n-1}(x^{n-1})\tilde{\pi}^{n}(x^{n})}\cdot\frac{|\det(J_{T^{n}}(x^{n-1}))|}{|\det(J_{T^{n}}(x^{n-1}_{*}))|}\right].$ (31)
Proof.
Let $S:=\{(z^{n-1},z^{n})\in\mathbb{R}^{d}\times\mathbb{R}^{d}\textrm{ such that }T^{n}(z^{n-1})=z^{n}\}$. Recall the definition of the extended measures ${\mathbb{P}}^{n-1}$ and ${\mathbb{Q}}^{n}$:
${\mathbb{P}}^{n-1}(\mathrm{d}z^{n-1},\mathrm{d}z^{n})$ $=\pi^{n-1}(\mathrm{d}z^{n-1})P^{n-1}(z^{n-1},\mathrm{d}z^{n}),$
${\mathbb{Q}}^{n}(\mathrm{d}z^{n-1},\mathrm{d}z^{n})$ $=\pi^{n}(\mathrm{d}z^{n})Q^{n}(z^{n},\mathrm{d}z^{n-1}).$
Moreover for $(z^{n-1},z^{n})\in S$,
${\mathbb{P}}^{n-1}(\mathrm{d}z^{n-1},\mathrm{d}z^{n})=(T^{n}\#\pi^{n-1})(\mathrm{d}z^{n})Q^{n}(z^{n},\mathrm{d}z^{n-1}).$
Therefore
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(z^{n-1},z^{n})=\frac{\pi^{n}(\mathrm{d}z^{n})}{(T^{n}\#\pi^{n-1})(\mathrm{d}z^{n})}\\ =\frac{\tilde{\pi}^{n}(z^{n})}{\tilde{\pi}^{n-1}((T^{n})^{-1}(z^{n}))|\det(J_{(T^{n})^{-1}}(z^{n}))|}=\frac{\tilde{\pi}^{n}(z^{n})|\det(J_{T^{n}}(z^{n-1}))|}{\tilde{\pi}^{n-1}(z^{n-1})}$ (32)
which justifies the identity for the work. Applying this at $(z^{n-1},z^{n})=(x^{n-1},x^{n}_{*})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1},x^{n}_{*})=\frac{\tilde{\pi}^{n}(x^{n}_{*})|\det(J_{T^{n}}(x^{n-1}))|}{\tilde{\pi}^{n-1}(x^{n-1})}.$ (33)
Similarly, applying Equation˜32 at $(z^{n-1},z^{n})=(x^{n-1}_{*},x^{n})\in S$ gives
$\frac{\mathrm{d}{\mathbb{Q}}^{n}}{\mathrm{d}{\mathbb{P}}^{n-1}}(x^{n-1}_{*},x^{n})=\frac{\tilde{\pi}^{n}(x^{n})|\det(J_{T^{n}}(x^{n-1}_{*}))|}{\tilde{\pi}^{n-1}(x^{n-1}_{*})}.$ (34)
Together Equation˜33 and Equation˜34 establish the proposition. ∎
C.1.2 Training
Since APT provides approximate samples from both the target and reference densities for each flow, it enables a range of training objectives. Some choices include maximum likelihood estimation (MLE) (equivalent to forward KL), reverse KL, and symmetric KL (SKL), which averages the two. Each has trade-offs: forward KL promotes mode-covering, reverse KL is more mode-seeking, and SKL balances both. APT’s parallel structure makes SKL particularly effective by providing access to samples at each intermediate annealing distribution, a feature many other methods lack. For example, sequential Monte Carlo (SMC)-based approaches such as FAB (Midgley et al., 2023) and CRAFT (Matthews et al., 2022) rely on samples from only one side, limiting their choice of loss functions.
Additionally, one can explore loss functions based on APT’s rejection rates, such as the analytic round trip rate, see Proposition˜2. Since higher round trip rates indicate more efficient mixing, optimizing for this metric improves sampling performance. Empirically, we found that using SKL yielded the most stable and robust results across different settings. This loss was therefore used in our final experiments. However, we leave it to future work to study other possible losses in more detail.
There are also possible variants to the training pipeline. In each case, we initialize normalizing flows to the identity transformation. One option is to run the APT algorithm with the current flows. After every (or several) steps of the APT algorithm, the current samples can be used to update the flow parameters. Since the flows are initialized at the identity transformation, the initial sampling of APT behaves similarly to PT. A second possible training pipeline is to instead use PT directly to generate a large batch of samples, and then use this batch of samples to update the flows. The latter approach is more stable and is thus what we employed in the final experiments. We provide more details in Appendix˜D. We note that the first training pipeline has the potential to allow for better exploration, and we therefore leave a more detailed exploration of it to future work.
C.2 Further Details on Control Accelerated PT
At the limit $K\rightarrow\infty$, following (Berner et al., 2025, Lemma B.7), and by the controlled Crooks fluctuation theorem (Vaikuntanathan & Jarzynski, 2008; Vargas et al., 2024), we arrive at the generalised work functional
$W^{n}_{\infty}(x)$ $:=\Delta F_{n}-\log\frac{\mathrm{d}{\mathbb{Q}}^{n}_{\infty}}{\mathrm{d}{\mathbb{P}}^{n-1}_{\infty}}$
$=\int_{0}^{1}-\nabla\cdot b^{n}_{s}(x_{s})+\nabla U^{n}_{s}(x_{s})\cdot b^{n}_{s}(x_{s})+\partial_{s}U^{n}_{s}(x_{s})\mathrm{d}s,$
inducing the corresponding continuous processes $(X_{s})_{s\in[0,1]}$, $(X_{s}^{\prime})_{s\in[0,1]}$ for ${\mathbb{P}}^{n-1}_{\infty}$, ${\mathbb{Q}}^{n}_{\infty}$ respectively, that is for the forward process
$X_{0}\sim\pi^{n-1},\quad\mathrm{d}X_{s}=-(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X_{s})\mathrm{d}s+b_{s}^{n}(X_{s})\mathrm{d}s+\sigma^{n}_{s}\sqrt{2}\mathrm{d}\accentset{\rightharpoonup}{B}_{s},$
and for the backwards process
$X^{\prime}_{1}\sim\pi^{n},\quad\mathrm{d}X^{\prime}_{s}=(\sigma^{n}_{s})^{2}\nabla U^{n}_{s}(X^{\prime}_{s})\mathrm{d}s+b_{s}^{n}(X^{\prime}_{s})\mathrm{d}s+\sigma_{s}^{n}\sqrt{2}\mathrm{d}\accentset{\leftharpoonup}{B}_{s}.$
The training objective is
$\mathcal{L}(b_{s},\phi_{s},\sigma_{s})=\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{K},{\mathbb{Q}}^{n}_{K})\xrightarrow[K\to\infty]{}\sum_{n=1}^{N}\mathrm{SKL}({\mathbb{P}}^{n-1}_{\infty},{\mathbb{Q}}^{n}_{\infty}).$
Note that the discrete version discussed in the main text is not the only discretisation choice. Other options (Albergo & Vanden-Eijnden, 2024; Máté & Fleuret, 2023) may also be applied.
C.3 Further Details on Diff-APT
Annealing path
We use the following Variance-Preserving (VP) diffusion process
$dY_{s}=-\beta_{s}Y_{s}\mathrm{d}s+\sqrt{2\beta_{s}}\mathrm{d}W_{s},\quad Y_{0}\sim\pi,$
with the choice of schedule $\beta_{s}=\frac{1}{2(1-s)}$ to define the path of distributions $(\pi^{\text{VP}}_{s})_{s\in(0,1]}$ by $Y_{s}\sim\pi_{1-s}^{\text{VP}}$. Due to the singularity at $s=1$, $\pi_{0}^{\text{VP}}$ is not defined by the path, but we can define this point to be a standard Gaussian as the path converges to this distribution in the limit as $s$ approaches 1 in order to define the full annealing path $(\pi_{s}^{\text{VP}})_{s\in[0,1]}$.
Accelerators
For a given annealing schedule $0=s_{0}<\ldots<s_{N}=1$, we construct $P_{k}^{n-1},Q_{k-1}^{n}$ through the linear discretisation of the time-reversal SDE on the time interval $s_{n-1},s_{n}]$ with step size $\delta_{n}=(s_{n}-s_{n-1})/K$ and interpolating times $s_{n,k}=s_{n-1}+k\delta_{n}$.
In particular, we take $Q_{k-1}^{n}(x_{k},dx_{k-1})$ to be ${\mathcal{N}}(\sqrt{1-\alpha_{n,k-1}}x_{k},\alpha_{n,k-1}\mathrm{I})$ where $\alpha_{n,k-1}=1-\exp(-2\int^{1-s_{n,k-1}}_{1-s_{n,k}}\gamma_{s}\mathrm{d}s)$ which is the closed-form kernel transporting $\pi_{s_{n,k}}^{\text{VP}}$ to $\pi_{s_{n,k-1}}^{\text{VP}}$. Furthermore, we take $P_{k}^{n-1}(x_{k-1},dx_{k})$ to be the exponential integrator given by ${\mathcal{N}}(\mu_{n,k-1}(x_{k-1}),\alpha_{n,k-1}\mathrm{I})$ where
$\mu_{n,k-1}(x)=\sqrt{1-\alpha_{n,k-1}}x+2(1-\sqrt{1-\alpha_{n,k-1}})(x+\nabla\log\pi_{s_{n,k-1}}^{\text{VP}}(x)).$
Network parametrisation
[…]
$\log g_{\theta}(x,s)$ $=[r_{\theta}(1)-r_{\theta}(s)][r_{\theta}(s)-r_{\theta}(0)]\langle N_{\theta}(x,s),x\rangle$
$+$ $[1+r_{\theta}(1)-r_{\theta}(s)]\log g_{0}(\sqrt{s}x),$
where $g_{1}(x)\propto\pi(x)N(x;0,\mathrm{I})$. Here, $r$ is a scalar-valued neural network and $N$ is a vector-valued function in $\mathbb{R}^{d}$. We also note that $\pi^{\theta}_{1}(x)\propto\pi$, hence $\pi^{\theta}_{s}$ serves as a valid annealing path between $N(0,\mathrm{I})$ and $\pi$.
Appendix D Experimental Details
D.1 Target Distributions
GMM-$d$
We take the 40-mode Gaussian mixture model (GMM-2) in 2 dimensions from Midgley et al. (2023) where to extend this distribution to higher dimensions $d$, we extend the means with zero padding to a vector in $\mathbb{R}^{d}$ and keep the covariances as the identity matrix but now within $\mathbb{R}^{d}$. This helps to disentangle the effect of multi-modality from the effect of dimensionality on performance as we essentially fix the structure of the modes across different values of $d$. Following previous work (Akhound-Sadegh et al., 2024; Phillips et al., 2024), we also scale the distribution GMM-$d$ by a factor of 40 to ensure the modes are contained within the range $1,1]^{d}$ for our experiments with APT, where we also use the same scaling for the PT baseline to ensure a fair comparison.
DW-4
We take the DW-4 target from Köhler et al. (2020) which describes the energy landscape for a toy system of 4 particles $\{x_{1},x_{2},x_{3},x_{4}\}$ and $x_{i}\in\mathbb{R}^{2}$ given by
$\pi(x)\propto\exp\left(-\frac{1}{2\tau}\sum_{i\neq j}a(d_{ij}-d_{0})+b(d_{ij}-d_{0})^{2}+c(d_{ij}-d_{0})^{4}\right),$
[…]
MW-32
We take the ManyWell-32 target from Midgley et al. (2023) formed from concatenating 16 copies of the 2-dimensional distribution
$\hat{\pi}(x_{1},x_{2})\propto\exp\left(-x_{1}^{4}+6x_{1}^{2}+\frac{1}{2}x_{1}-\frac{1}{2}x_{2}^{2}\right),$
i.e. the distribution $\pi(x)=\prod_{i=1}^{16}\hat{\pi}(x_{2i-1},x_{2i})$ where $x\in\mathbb{R}^{32}$. Each copy of $\hat{\pi}$ has 2 modes and hence $\pi$ contains in total $2^{16}$ modes.
Alanine Dipeptide
This is a small molecule with 22 atoms, each of which has 3 dimensions. The target energy is defined with the amber14/protein.ff14SB forcefield in vacuum using the DMFF library in JAX (Wang et al., 2023).
D.2 Network and Training Details
NF-APT
For GMM-$d$, we use 20 RealNVP layers where we employ a 2-layer MLP with 128 hidden units for the scale and translation functions (Dinh et al., 2017). We initialize the flow at the identity transformation. We use the Adam optimizer with a learning rate of 1e-3, perform gradient clipping with norm $1$ and employ EMA with decay parameter $0.99$.
We use the same training pipeline as for CMCD-APT. See further details in the CMCD-APT description below. Note that we also employ the SKL loss for training and the linear annealing path with a standard Gaussian as the reference distribution.
CMCD-APT
For both GMM-$d$ and MW-32, we use a 4-layer MLP with 512 hidden units, and for DW-4, we use a 4-layer EGNN (Satorras et al., 2021) with 64 hidden units. Recall that in CMCD-APT, we use the linear path, $U^{\beta}(x)=(1-\beta)\log\eta(x)+\beta U(x)$, which linearly interpolates between reference and target in log-space. Therefore, our network is conditional on the $\beta$. We optimise the MLP by Adam with a learning rate of 1e-3 and EGNN with a learning rate of 1e-4. We additionally use gradient clipping with norm 1 for stability.
Furthermore, our training pipeline following the 3 stages outlined below:
  • Tuning PT: We define $\{\beta_{n}\}_{n=1}^{N}$ uniformly from 0 to 1. Then, we run PT for 600 steps, remove the first 100 steps as burn-in, and take the last 500 steps to calculate the rejection rate between adjacent chains. We then tune the value of $\beta_{n}$ to ensure the rejection rate even, according to Syed et al. (2021). We repeat this process 10 times to ensure $\{\beta_{n}\}_{n=1}^{N}$ is stable.
  • Collecting data and training: We use PT with the tuned schedule to collect data. For experiments with 5 chains, we run 200K steps to collect 200K samples for each chain. For experiments with 10 and 30 chains, we run 65536 steps to collect data. We then train CMCD for 100,000 iterations with a batch size of 512. In each batch, we randomly select the chain index and samples according to the chain to train CMCD using SKL. We repeat this step twice to ensure the network is well-trained until convergence.
  • Testing: We follow Algorithm˜1, running CMCD-APT for 100K steps, and calculate the round trip.
Diff-APT
For GMM-$d$, we use a 3-layer MLP with 128 hidden units, for MW-32, we use a 3-layer MLP with 256 hidden units and for DW-4, we use a 3 layer EGNN with 128 hidden units. For all models, we use a learning rate of 5e-4, gradient clipping with norm 1 for stability and EMA with decay parameter of 0.99. For training, we follow the same pipeline as CMCD-APT, but we only retain samples at the target chain in order to optimise the standard score matching objective. At sampling time, we tune the annealing schedule by running Diff-APT for 1,100 steps, discarding the first 100 samples as burn-in and using the last 1,000 steps to calculate rejection rates to apply the schedule tuning algorithm from Syed et al. (2021). We then repeat this 10 times where we initialise from the uniform schedule.
PT
For all experiments with PT, we use the linear path with a standard Gaussian as our reference distribution. We tune the annealing schedule in the same manner as with Diff-APT to ensure comparisons with a strong baseline.
D.3 Further Details on Comparison of Acceleration Methods
For both training and testing for all methods, we take a single step of HMC with step size of $0.03$ and 5 leapfrog steps as our local exploration step across each annealing chain. We note that while we could have improved performance by tuning the step size for each annealing distribution, we keep this fixed to disentangle the effect of local exploration from our communication steps.
Compute-normalised round trips
The compute-normalised round trips, as reported in Table 1, is computed by dividing the original round trips by the number of potential evaluations that a single “machine” is required to implement within a parallelised implementation of PT/APT - i.e. the number of potential evaluations required by the computation of $W^{n}_{K}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0:K})$ (or equivalently $W_{K}^{n}(\accentset{\leftharpoonup}{X}_{t,0:K}^{n}$) for a single $n$ and $t$. Similarly, we count the number of neural calls in the corresponding manner.
  • NF-APT: We need 2 potential evaluations for $\pi^{n-1}(\accentset{\rightharpoonup}{X}^{n-1}_{t,0})$ and $\pi^{n}(\accentset{\rightharpoonup}{X}^{n-1}_{t,1})$ and single network evaluation.
  • CMCD-APT: For $K>0$, we need to calculate the potential and score [^3] We count this as a single potential evaluation. of $\pi^{n-1}$ at $\accentset{\rightharpoonup}{X}_{t,0}^{n-1}$ and $\pi^{n}$ at $\accentset{\rightharpoonup}{X}^{n-1}_{t,K}$. We also need to compute the score of $U^{n}_{s_{k}}$ at $\accentset{\rightharpoonup}{X}_{t,k}^{n-1}$ for $k=1,\ldots,K-1$. We note that we can reuse all of the above score evaluations for both the forward and reverse transition kernels of $P_{k}^{n-1}$ and $Q_{k-1}^{n}$. In total, this requires $K+1$ potential and network evaluations.
  • Diff-APT: For $K>0$, we require $K+1$ potential and network evaluations following the same logic as for CMCD. For $K=0$, we require 2 potential and network evaluations for $\pi^{n-1}(X^{n-1})$ and $\pi^{n}(X^{n-1})$ as we parametrise our annealing path in terms of the target distribution $\pi$ and a neural network.
  • PT: We need 2 potential evaluations for $\pi^{n-1}(X^{n-1})$ and $\pi^{n}(X^{n-1})$ and we do not require any network evaluation.
D.4 Further Details on Scaling with Dimension
For all methods, we take a single step of HMC with step size of $0.03$ and 5 leapfrog steps as our local exploration step across each annealing chain and take 100,000 samples. Additionally, we report the (compute-normalised) round trip rate which involves dividing the (compute-normalised) round trips by the number of sampling steps. We note that we use the same methodology as above for computing compute-normalised round trips.
D.5 Further Details on Free Energy Estimator
For CMCD-APT, we take a single step of HMC with step size 0.22 and 5 leapfrog steps for our local exploration step across each annealing chain for both DW-4 and MW-32. For Diff-APT, we take two steps of HMC with 5 leapfrog steps and step size of 0.2 and 0.22 respectively for DW-4 and MW-32 for our local exploration step across each annealing chain.
For all methods, we generate 100,000 samples at the target distribution. For each estimate of $\Delta F$, we subsample 1,000 samples uniformly without replacement from the 100,000 to compute the APT free energy estimator. This is then repeated 30 times for each method.
We take the ground truth free energy of DW-4 from estimating $\Delta F$ with the APT estimator using 100,000 samples from PT with 60 chains (after tuning). We take the ground truth free energy of MW-32 from Midgley et al. (2023) which calculates the normalising constant of a single copy of $\hat{\pi}$ numerically allowing for the trivial computation of the overall normalising constant.
D.6 Further Details on Comparing APT with Neural Samplers
For CMCD, we take our CMCD-APT model on DW-4 with 30 chains and $K=1,2,5$ from Section 6.3, and instead of sampling from these models using APT, we collect 5,000 independent samples from our reference distribution to which we apply our learned kernels $\prod_{n=1}^{N}\prod_{k=1}^{K}P_{k}^{n-1}$ which we recall are explicitly trained to transport samples from the reference to the target distribution. With our final samples, we then plot the histogram of $d_{ij}$ values for $i\neq j$.
For Diffusion, we follow the same procedure but we tune our annealing schedule with the same schedule tuning algorithm from Syed et al. (2021) as mentioned above (this step is not required for CMCD as the annealing schedule is required to be fixed during training) before we collect samples by applying $P_{k}^{n-1}$.
For CMCD-APT and Diff-APT, we simply sample from each method to generate 5,000 samples at the target distribution before plotting the histogram of $d_{ij}$ values.
D.7 Further Details on Alanine Dipeptide
We run CMCD-APT using 5 chains and $K=1,2,5$. For the local move, we adopt HMC with a dynamic step size: when the acceptance rate is larger than 0.9, we increase the step size by 1.2; if the acceptance rate is smaller than 0.8, we divide the step size by 1.2. Other settings are the same as those of other targets.
D.8 License
Our implementation is based on the following codebases:
D.9 Computing Resources
The experiments conducted in this paper are not resource-intensive. We use a mixture of 24GB GTX 3090 and 80GB A100 GPU, but all experiments can be conducted on a single 80GB A100 GPU.
Appendix E Additional Experimental Results
E.1 Visualisation of ManyWell-32
In Figures˜5 and 6, we visualise ManyWell-32 samples generated by 1,000 consecutive CMCD-APT and Diff-APT steps with $N=5,10,30$ and $K=1,5$ from ManyWell-32 via marginal projections over the first four dimensions. We also show 1,000 independent ground truth samples in Figure˜7 for comparison. We choose to report only 1,000 steps to illustrate how fast our sampler mixes. As we can see, the mode weights become more accurate as we increase $N$ and $K$.
(a) $N=5,K=1$.
(b) $N=5,K=5$.
(c) $N=10,K=1$.
(d) $N=10,K=5$.
(e) $N=30,K=1$.
(f) $N=30,K=5$.
Figure 5: Visualisation of ManyWell-32 samples generated by 1,000 consecutive CMCD-APT steps.
(a) $N=5,K=1$.
(b) $N=5,K=5$.
(c) $N=10,K=1$.
(d) $N=10,K=5$.
(e) $N=30,K=1$.
(f) $N=30,K=5$.
Figure 6: ManyWell-32 samples generated by 1,000 consecutive Diff-APT steps.
Figure 7: 1,000 independent ground truth samples from ManyWell-32
E.2 Comparison of Acceleration Methods for ManyWell-32 and DW-4
We take our trained models from Section 6.3 and provide the same comparison with PT as in Table 1 for ManyWell-32 and DW-4 below.
Table 3: PT versus APT with different acceleration methods, targeting ManyWell-32 in 32 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per potential evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration, and $\Lambda$ estimated using $N=30$ chains.
# Chain $N=5$ $N=10$ $N=30$
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
CMCD-APT ($K=1$) 2 4.384 1154 577.0 2802 1401.0 4729 2364.5
CMCD-APT ($K=2$) 3 3.827 1587 529.0 3640 1213.3 5544 1848.0
CMCD-APT ($K=5$) 6 3.148 2878 479.7 4790 798.3 6678 1113.0
Diff-APT ($K=1$) 2 6.663 425 212.5 2402 1201 4398 2199
Diff-APT ($K=2$) 3 5.225 1387 462.3 4022 1340.7 5894 1964.7
Diff-APT ($K=5$) 6 3.94 3627 604.5 5704 950.7 7634 1272.3
Diff-PT ($K=0$) 2 7.423 251 125.5 1561 780.5 3440 1720
PT 0 5.475 550 275 1879 939.5 3733 1866.5
Table 4: PT versus APT with different acceleration methods, targeting DW-4 in 10 dimensions and standard Gaussian reference using $N=5,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per potential evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration, and $\Lambda$ estimated using $N=30$ chains.
# Chain $N=5$ $N=10$ $N=30$
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
CMCD-APT ($K=1$) 2 3.173 3020 1510.0 6407 3203.5 9456 4728.0
CMCD-APT ($K=2$) 3 2.671 4239 1413.0 7549 2516.3 10538 3512.7
CMCD-APT ($K=5$) 6 2.107 6971 1161.8 9808 1634.7 12634 2105.7
Diff-APT ($K=1$) 2 4.565 4331 2165.5 7397 3698.5 7729 3864.5
Diff-APT ($K=2$) 3 3.810 7187 2395.7 10176 3392 9176 3058.7
Diff-APT ($K=5$) 6 4.358 12456 2076 12740 2123.3 8104 1350.7
Diff-PT ($K=0$) 2 4.739 2962 1481 5862 2921 7067 3533.5
PT 0 4.016 2329 1164.5 5128 2564 7610 3805